### v1

# SelfAI: Building a Self-Training AI System with LLM Agents

Abstract
Recent work on autonomous scientific discovery has leveraged LLM-based agents to integrate problem specification, experiment planning, and execution into end-to-end systems. However, these frameworks are often confined to narrow application domains, offer limited real-time interaction with researchers, and lack principled mechanisms for determining when to halt exploration, resulting in inefficiencies, reproducibility challenges, and under-utilized human expertise. To address these gaps, we propose SelfAI, a general multi-agent platform that combines a User Agent for translating high-level research objectives into standardized experimental configurations, a Cognitive Agent powered by LLMs with optimal stopping criteria to iteratively refine hyperparameter searches, and an Experiment Manager responsible for orchestrating parallel, fault-tolerant training workflows across heterogeneous hardware while maintaining a structured knowledge base for continuous feedback. We further introduce two novel evaluation metrics, Score and $\text{AUP}_{D}$, to quantify discovery efficiency and search diversity. Across regression, NLP, computer vision, scientific computing, medical imaging, and drug discovery benchmarks, SelfAI consistently achieves strong performance and reduces redundant trials compared to classical Bayesian optimization and LLM-based baselines, while enabling seamless interaction with human researchers.
1 Main
In recent years, large language models (LLMs) [1, 2, 3] have fundamentally reshaped the landscape of AI-driven research. Advances in reasoning [4, 5], multimodal understanding [6, 7], and autonomous tool use [8, 9, 10, 11, 12, 13] have positioned LLMs as central components for accelerating scientific workflows.
Early LLM-based scientific research demonstrated that LLMs can extract actionable scientific knowledge to guide experiments [14, 15, 16, 17] or answer professional questions [18, 19, 20]. As reasoning and tool-interaction capabilities matured, research expanded to cross-stage planning [21, 22, 23], hypothesis and idea generation [24, 25, 26, 27], multimodal knowledge integration [28], and conducting experiments [17, 29]. Meanwhile, LLMs have driven the development of tasks such as reaction prediction [30, 31, 32], material discovery [33, 34, 18, 35], chemical synthesis design and molecular property optimization [36, 29, 37, 30], and biomedicine research [17, 38, 39, 40, 41]. Building on these advances, a new generation of scientific discovery systems has emerged, spanning AI scientific assistants across the full spectrum of scientific workflows [18, 42, 16, 43], automated scientific discovery systems (ASDSs) specialized for code generation and experiment design [36, 44, 29], and increasingly capable research agents that incorporate automated debugging, iterative idea refinement, and scientific writing [26, 45, 46]. Along this evolutionary trajectory, LLM-centered ASDS have begun to integrate inference, execution, and feedback mechanisms into unified closed-loop pipelines, achieving varying degrees of autonomy across scientific domains.
Despite rapid progress, existing ASDS frameworks still exhibit fundamental limitations. Many systems [44, 45, 24, 47] (such as AIRA and Scientist-V2) excel at translating research intent into executable code or experimental procedures, thereby validating the “executability” of ASDS and primarily relying on final success rate (medal rate) or best performance as metrics. More autonomous platforms like MLGym [48] and MLAgentBench [49] enable system-level control, allowing agents to execute end-to-end benchmarks, collect results, and conduct standardized evaluations. Moreover, while recent LLM-driven optimization methods [50, 51, 52, 53] such as LLM4EO showcase that LLMs can infer evolutionary tendencies and synthesize more effective operators, they operate at the level of local modification rules rather than scientific reasoning in trajectories. Overall, these approaches do not explicitly enhance scientific reasoning during experimentation, such as identifying optimal stopping points, discovering efficient trial trajectories, or assessing the structural quality of reasoning.
Building reasoning-centric ASDS therefore requires a shift from execution-oriented autonomy toward cognitive autonomy: systems must not only carry out experiments but also analyze, reflect, and revise their reasoning throughout the experimental process. However, current ASDS systems provide little support for these cognitive-layer behaviors. The absence of mechanisms for trajectory-level reasoning assessment and LLM-driven decision-making limits both efficiency and accuracy, especially in settings with large search spaces, stochastic outcomes, or complex interdependencies among hypotheses, resulting in inefficient exploration, elevated computational costs, and limited real-time adaptability. These limitations highlight a fundamental need for ASDS architectures that incorporate reasoning-driven search, adaptive learning mechanisms, and structural evaluation of scientific trajectories.
Figure 1: SelfAI Framework for Automated Scientific Experimentation. a, Holistic architecture of the multi-agent system, which transforms various experiments in the research process into a structured workflow. b, User intentions, comprising ideas and experiment schemes, are transformed into structured configurations via a predefined prompt. These inputs are processed through successive stages: hypothesis generation, strategic planning, trial execution, and result collection. c, Performance distribution across 11 tasks demonstrates the framework’s ability, when powered by GPT-4o-mini, to prioritize high-performance regions without sacrificing exploration. d, Trial counts for each solver shown in c, accompanied by quantile lines, density distributions, and performance variability across the global and two evaluation regions. Higher values in low-performance regions promote rapid escape, while lower values in high-performance regions enable localized refinement.
To address these issues, we propose SelfAI, a unified multi-agent self-training pipeline for continuous adaptability and transparent collaboration. As shown in Fig. 1, SelfAI forms a closed loop that integrates user intent, cognitive reasoning, and experimental orchestration. Upstream, a User Agent converts high-level objectives and exploratory questions into standardized experiment configurations. These configurations enable the Cognitive Agent to analyze performance metrics, conduct reasoning over historical outcomes, and adjust the search trajectory, driving gradual improvement [54, 55]. A resource-aware Experiment Manager ensures efficient execution by handling resource scheduling, environment provisioning, adaptive checkpointing, and comprehensive experiment logging. During high-level parameter optimization, the Cognitive Agent and Experiment Manager jointly estimate the potential benefit of continued training and apply an optimal-stopping criterion to terminate unpromising trials. To evaluate scientific reasoning quality, we introduce two complementary evaluation metrics, Score (Discovery Efficiency) and $\text{AUP}_{D}$ (Area Under the Performance Diversity). Score aggregates, across tasks, the normalized improvement over the search space together with penalties for discovering good configurations late and for stopping far from the best-found point. $\text{AUP}_{D}$ explicitly encodes how broadly a solver explores the search space by summarizing the performance-diversity tradeoffs across the entire trajectory, enabling detailed analysis of exploration behavior and stopping decisions in long-term searches. Together, these metrics enable quantitative assessment of exploration behavior, reasoning structure, and stopping decisions in long-term autonomous experiments.
In summary, SelfAI provides a continuously self-training system that autonomously iterates through cycles of intent-centric understanding, trajectory-aware reasoning, and adaptive strategic planning, enabling sustained performance gains with minimal human intervention. Through checkpoint tolerance, zero-code parallelization, and dynamic resource management, SelfAI substantially reduces redundant computation, improves resource efficiency, and accelerates search convergence in long-horizon scientific workflows. In addition, SelfAI integrates seamlessly with modern MLOps pipelines [56, 57, 58], facilitating collaborative experiment coordination, end-to-end traceability of model evolution, and robust privacy protection for sensitive configurations and user data. Collectively, these capabilities make SelfAI a practical, scalable, and operationally mature platform for rapid scientific discovery. The key contributions of this paper are given as follows:
  1. 1.
    Autonomous Scientific Discovery System: We design a cohesive framework of specialized LLM agents, including the User Agent, Cognitive Agent Framework, and Experiment Manager, which collaboratively translate high-level research objectives into iteratively refined experimental workflows, including trajectory analysis, hypothesis generation, and best stopping decision-making.
  2. 2.
    Novel Evaluation Metrics and Optimal Stopping: We propose two new metrics, Score (discovery efficiency) and $\text{AUP}_{D}$ (area under the performance–diversity curve), to jointly quantify the efficiency and diversity of hyperparameter exploration. An embedded optimal stopping criterion automatically terminates unpromising trials, effectively balancing exploration and exploitation while minimizing computational overhead.
  3. 3.
    Comprehensive Validation of LLM-based Reasoning in Scientific Discovery: We evaluate SelfAI across diverse tasks, including regression, sentiment analysis, computer vision, medical imaging, and drug discovery. The results show that SelfAI effectively improves upon excellent baseline models in both results exploration and performance.
2 Performance of SelfAI on Benchmark Data
Figure 2: Scores among all solvers across different tasks to measure the best stopping criterion.
Figure 3: Diverse Metrics ($\text{AUP}_{D}$) among all solvers across different tasks to evaluate trajectory diversity.
Scientific discovery is driven by diverse physical environments and user intents. To effectively reason across diverse scientific discovery tasks, our approach integrates multiple LLM agents and a suite of tools: User Agent, Cognitive Framework, and Experiment Management, addressing the complex reasoning and planning problem of scientific discovery. In our experiments, we evaluate the SelfAI across 6 domains and 12 tasks. The compared methods contain different suites of LLMs, such as OpenAI-o3 [1], Llama3.3 [59], Qwen2.5 [2], and DeepSeek-R1 [3]. We used grid search and the Tree Parzen Estimator (TPE) optimizer [60, 61] (referred to as BS), which applies a modified Bayesian inference method using a spiral search. All agents are tested on four Nvidia A100 GPUs. A public implementation of SelfAI will be available on the project website. We also provide reasoning processes for reproducibility.
[…]
Benchmarks. We collect benchmarks composed of multiple tasks of 6 primary categories, designed to simulate challenging scenarios in scientific discovery. The benchmark requires the solver to engage in scientific reasoning by analyzing experiments, interpreting observations, and adapting strategies, uncovering novel patterns, functional relationships, or optimal experimental configurations. Compared with hyperparameter search, these processes reflect the inductive reasoning essential to real-world scientific inquiry. Therefore, this benchmark assesses the model’s reasoning ability, not just its numerical optimization capabilities. All benchmark data is collected through actual experimental runs, and some results are reported in published literature, spanning both discrete and continuous hyperparameter search spaces, from low-dimensional to high-dimensional settings (see Appendix B for more details).
Existing methods [24, 45, 49] explore final performance or medal rates using reward functions and prompts specifically designed for innovative idea generation, code generation, and pass@K strategies. In the line of LLM-driven optimization, LLM4EO [50] uses LLMs exclusively to synthesize and evolve evolutionary operators that determine which parts of a candidate solution should be modified during the search. The LLM is used to generate and update operators to shape the local search behavior of the evolutionary algorithm. These implementations do not focus on scientific reasoning across trials during the exploration process, making direct comparisons with our benchmark infeasible. Consequently, we used the readily accessible GPT4-o3-mini model and compared SelfAI with LLM and LLM-ES methods on 11 small/medium-sized tasks across different domains in Fig. 1c. The LLM and LLM-ES methods are fine-tuned by using the structured prompts in Supplementary Section D. The LLM solver uses a search prompt that recommends new trials based on prior results, while LLM-ES augments this with an early-stopping module that decides whether to terminate exploration. Fig. 1c presents that our SelfAI outperforms LLM and LLM-ES in all four metrics, validating SelfAI’s effectiveness in integrating domain knowledge to accelerate scientific discovery. Furthermore, we focus on the reasoning capability of the original LLM model because it is difficult to be fair when comparing different methods. Fig. 1d shows that SelfAI promptly focuses on exploiting higher-performance regions while maintaining a balanced distribution for effective exploration. Figs. 2 and 3 summarize the benchmark performance based on these two metrics. The benchmarks show that SelfAI consistently achieves excellent results across various domains and tasks for different LLM models. Detailed comparisons covering Score, $\text{AUP}_{D}$, Best-Time, Stop-Time, and Hit Rate are provided in Supplementary Table 3.
Failure cases. Several key failure modes of LLM-based solvers are shown in Supplementary Section B.1. First, the hit rate for identifying the best result varies substantially across models, with some solvers (particularly DeepSeek-R1 series) failing to reach the global optimum due to premature stopping. Second, limited context windows cause incomplete reasoning over long trajectories, preventing solvers from fully incorporating earlier experimental signals. Third, computational fragility leads to non-monotonic performance under seemingly minor perturbations in reasoning strategy. These findings indicate that while SelfAI improves overall efficiency, reasoning-driven exploration remains sensitive to solver stability and stopping behavior.
2.1 Machine Learning
Boston house pricing prediction. We evaluate SelfAI on the Boston housing price prediction task [62] using a random forest regression model. A total of 162 trials were conducted across a search space defined by five hyperparameters: i.e., $\text{n-estimators}=[100,200,300]$, $\text{max-depth}=[\text{None},10,20]$, $\text{min-samples-split}=[2,5,10]$, min-samples-leaf=[1,2,5] and $\text{max-features}=[\text{``sqrt'',``log2''}]$). As shown in Figs. 2 and 3 (and Supplementary Table 4), GPT-4o-mini achieves by far the highest Score (0.9811) and the 1st ranking, indicating outstanding optimization efficiency. Within the DeepSeek-R1 family, the 32B and 7B variants secure 2nd and 3rd places respectively, demonstrating the best balance between rapid convergence (low Best-Time) and reasonable search diversity among open-source models. In contrast, DeepSeek-R1-70B and Llama3.3-70B exhibited consistent exploration but lacked effective stopping criteria, leading to longer convergence times. Notably, nearly all models achieved an identical and high Best Result ( 0.841), indicating that while most can find a correct solution, they differ substantially in optimization efficiency and reliability.
Sentiment analysis. We perform experiments for sentiment analysis that focus on identifying opinions, emotions, and attitudes expressed in text. All experiments are found in [63] with standard experimental settings, where the pre-trained Word2Vec embeddings [64] are used as input features. An LSTM network is then employed to model sentence sequences, converting them into a multi-class classification problem. This setup serves to evaluate the agent’s reasoning and optimization capabilities in textual domains. As shown in Supplementary Table 5, GPT-4o-mini outperforms all other LLMs, achieving the highest Score (0.8824) by discovering the optimal configuration extremely early while maintaining reasonable search diversity. DeepSeek-R1 series and Qwen2.5 series exhibit significantly lower Scores, with many failing to converge quickly or occasionally missing the global optimum. The largest frontier model GPT-4o also underperforms markedly (Score 0.2745, rank 11), suggesting that scale alone is insufficient for effective iterative hyperparameter reasoning in recurrent neural architectures.
2.2 Scientific Computing for Image Completion
For scientific computing fields, we selected the tensor decomposition method [65]. The method is an important tool for high-dimensional data analysis and is crucial in applications such as data compression [66], computational acceleration [67], and multi-modal data fusion [28]. All LLMs are provided with identical mathematical knowledge about TW decomposition generated by GPT-4o. As shown in Supplementary Table 6, Qwen2.5 (7b and 14b) and GPT4-o3-mini perform well and rank highly. Qwen2.5-72b, on the other hand, may overly rely on general reasoning capabilities and fail to balance mathematical knowledge and reasoning capabilities. In the DeepSeek series, although the DeepSeek-R1-7b model can’t find the optimal solution, its search strategy was acceptable with a moderate Score, demonstrating its ability to perform mathematical reasoning. DeepSeek-R1-70b and GPT-4o, despite excellent search diversity and early stopping, received a mediocre score, reflecting that their exploration strategy was not well aligned with the optimization goal of tensor decomposition, possibly failing to find the optimal performance between data compression and computational efficiency.
2.3 Computer Vision
A. SIREN We employed SIREN (Sinusoidal Representation Networks) [68] to evaluate our framework on image segmentation and denoising tasks. Leveraging sine activations and coordinate-based inputs, SIREN excels at representing high-frequency signals through continuous implicit representations, making it widely used in scientific computing and physics-based problems. However, its performance is highly sensitive to hyperparameters (e.g., learning rate and regularization strength, etc.), requiring careful tuning per dataset. The unsupervised nature of SIREN further increases the risk of training instability or divergence with improper settings, making it a strong test case for SelfAI’s hyperparameter optimization capability.
Fig. 4 illustrates hyperparameter search trajectories for image segmentation using SIREN (see also Supplementary Figs. 8, 9 and 10), where the surface, smoothed from the original steep, multi-peak data, reveals distinct search behaviors. Given the same three initial points, the tree-structured Bayesian optimizer (BS) follows a spiral trajectory that underutilizes promising starting regions and fails to explore broadly, often converging to local minima. In contrast, LLM-based optimizers infer trends from evaluated points, incorporate causal reasoning, and explore more broadly to efficiently locate the global optimum. They also monitor progress and halt when improvements plateau, a capability absent in traditional methods.
Quantitative results (Supplementary Tables 78) reveal distinct task-dependent performance profiles. In segmentation, DeepSeek-R1-7B achieves the highest Score (0.693) and ranks first, followed by Qwen2.5-7B and GPT-4-o3-mini. Larger models such as Qwen2.5-72B and Llama-3.3-70B rank only mid-range, indicating that scale alone does not ensure effective optimization. In denoising, the landscape shifts: Qwen2.5-72B delivers the best performance (Score 0.761), while DeepSeek-R1-14B and DeepSeek-R1-32B jointly occupy second place. Notably, models that perform well in segmentation tasks do not necessarily perform well in denoising tasks, suggesting that the effectiveness of the solver largely depends on the degree of matching between its inference strategy and the task.
B. Image Classification We also benchmark two commonly used supervised learning methods in computer vision: Masked Autoencoder (MAE) [69] and ResNet [70]. These represent fundamentally different learning paradigms where hyperparameter optimization plays a crucial role in achieving the latest performance.
Mask Autoencoder (MAE) MAE introduces two key hyperparameters: training strategies (linear detection vs. fine-tuning) and masking rate, which are explored over a range: $[0.10,0.90]$ with an interval of 0.1. The masking rate directly affects the difficulty of the reconstruction task and the quality of the learning representation, while the training strategy determines how to adapt to the downstream task. In our experiments (see Supplementary Table 9 for more details), GPT-4o-mini achieved explicit performance with efficiency, demonstrating the extraordinary ability to identify the best mask configuration and training strategies. Qwen2.5-7B achieves early convergence and disciplined stopping. The DeepSeek-R1 series shows consistent scaling behavior: performance improves from 7B to 14B to 32B, but slightly drops at 70B. In traditional solvers, GS and BS lag significantly behind LLM-based solvers, highlighting the advantages of intelligent hyperparameter optimization methods. Apart from Qwen2.5-14b and DeepSeek-R1-7b, most solvers can find the optimal solution, but significant differences exist among the various methods in terms of search efficiency and convergence speed.
ResNet Family On the ImageNet ResNet hyperparameter benchmark, LLMs in the 7B–32B range demonstrate superior optimization efficacy. The top-performing solvers, such as Qwen2.5-14B (1st), DeepSeek-R1-32B (2nd), and Qwen2.5-7B with DeepSeek-R1-14B (tied 3rd), consistently identify the optimal ResNet configuration within one or two trials, halting immediately with near-perfect sample efficiency. In contrast, larger variants from the same model families drop to middle or lower rankings, consuming over half the search budget with notably higher exploration overhead. While GPT-series models, as large-scale counterparts, still achieve competitive results, the 7B–32B class exhibits a clear advantage in balancing accuracy and efficiency. For the architectural search on ImageNet, we explored key dimensions such as depth and bottleneck design. Qwen2.5-14B attains theoretical peak performance, with its 7B and 32B versions completing further exploration efficiently. These results reflect an ability to identify ideal network configurations with minimal wasted exploration.
LCBench We evaluate SelfAI on the standard AutoML benchmark [71] using Bayesian optimization over 2000 rounds of hyperparameter search, covering critical parameters including learning rate, batch size, network depth, and dropout rate. As shown in Supplementary Table 11, DeepSeek-R1-70B, Qwen2.5-14B, and Llama3.3-70B secure the top three positions in the ranking, with DeepSeek-R1-70B attaining the highest overall score and the lowest search diversity. Notably, leading models in the 7B–32B parameter range efficiently identify near-optimal hyperparameter configurations by evaluating only a minimal set of candidate solutions, achieving an order-of-magnitude improvement in stop-time compared to classical methods. These findings demonstrate diminishing returns with scaling, as larger models within the same series (e.g., Qwen2.5-72B) typically underperform comparable models of medium size, while the 7B model can achieve robust performance comparable to models several times its size.
Figure 4: Illustration of the optimized trajectory for the SIREN method for image segmentation. Green points are suggested points before reaching the optimal points. Red points are redundant suggestions when reaching out to the optimal points and failing to stop trials. The $\star$ is the optimal point. We show the serialization recommendations provided by LLM through the labeled numbers.
2.4 Medical Image Analysis
In medical image analysis, nnU-Net [72] is a landmark framework known for its strong generalization and automated design. Despite its adaptive network structure, preprocessing, and training strategies for various segmentation tasks, the exploration of novel architectures persists. To address this, nnU-Net-Revisited [73] establishes a comprehensive benchmark including 19 mainstream models (CNN-based [74, 75, 76], Transformer-based [77, 78], and Mamba-based [79, 80]), offering a solid basis for fair evaluation. Results are shown in Figs. 2-3 and Supplementary Tables 1213. On the BTCV dataset [81], GPT4-o3-mini achieves the highest Score (0.6875) and Rank 1, demonstrating strong optimization efficacy, while Qwen2.5-7B secures second place, outperforming all larger variants in its family. In nnU-Net hyperparameter tuning for the BraTS dataset [82], Qwen2.5-14B attains the top ranking (Score 0.4333), followed closely by GPT4-o3-mini, with Qwen2.5-32B and DeepSeek-R1-70B tying for third. Notably, the 7B and 72B models outperformed similar models of medium size (14B–32B), indicating that larger model size does not guarantee better optimization.
2.5 Imbalanced Node Classification
Graph neural networks (GNNs), as powerful tools for processing relational data, are playing an increasingly critical role in numerous cutting-edge scientific fields. In biomedicine, GNNs have become a core component of protein structure prediction models such as AlphaFold3 [39], enabling precise modeling of spatial interactions between amino acid residues and ushering in a new era in structural biology. In drug discovery, they are widely used for molecular property prediction [39, 83, 84], compound-target interaction identification [85, 86], and novel drug generation [87]. Similarly, GNNs demonstrate exceptional performance in engineering domains, including traffic prediction [88, 89], cybersecurity, and chip design [90, 91]. Given this broad utility and the sensitivity of GNN performance to hyperparameter settings, we evaluate GraphSAGE on the imbalanced Cora benchmark [92] under the setup of [93].
Figs. 2 and 3 represent that the GPT4-o3 series delivered near-optimal results, achieving the fastest discovery of optimal solutions and efficient stopping decision while maintaining high search diversity, underscoring its robust optimization capability in scientific computing scenarios. In contrast, Qwen2.5 models, (except 72b), easily delivered consistent but middling results across scales, while larger models generally exhibit late convergence and higher unnecessary exploration, reflecting potential inherent limitations in understanding the structural properties and optimization requirements of specialized domains such as protein interaction networks and molecular graph structures (see Supplementary Table 14 for more results). These results indicate that SelfAI can effectively address challenges in GNN hyperparameter optimization. The framework’s strong performance on graph learning tasks suggests its potential as a general-purpose optimizer for scientific domains, including protein engineering and drug development.
2.6 Drug Discovery
In drug discovery, accurate bioactivity prediction is vital for early-stage virtual screening and compound prioritization, enabling shorter development cycles and lower preclinical costs. Following the evaluation practices and modeling standards summarized by Korotcov et al. [94], this study employs the Chagas EP20 dataset [95], which measures the activity of compounds against Trypanosoma cruzi, a neglected but medically significant target in tropical disease research. The core challenge lies in learning an effective mapping from molecular representations, such as SMILES sequences [96] or graph-based encodings [97] to experimentally measured bioactivity [98]. We adopt SelfAI to adaptively refine performance, assess search trajectories, and determine principled stopping points for this task.
Experimental results (see Supplementary Table 15) reveal distinct performance patterns among solvers in hyperparameter optimization. Qwen2.5-7B and 14B secure first and second place, respectively, both achieving high Scores with minimal search diversity. GPT-4o ranks third, while DeepSeek-R1-7B and 70B tie for fourth. A key finding is the clear performance drop at larger scales; Qwen2.5-32B, Qwen2.5-72B, and DeepSeek-R1-32B rank in the bottom half, demonstrating that increased parameters do not guarantee better optimization. In contrast, traditional optimization methods (such as GS and BS) ranked poorly in this task, underscoring the clear advantage of large language model-based intelligent optimizers in bioactivity prediction. These results collectively demonstrate that in high-dimensional, noisy biochemical scenarios, language model-driven optimizers offer efficient convergence and a superior balance between exploration and exploitation.
Finally, the ranking heatmap (Fig. 7) and average performance analysis (Supplementary Table 3) reveal several consistent trends. GPT-4o-mini emerges as the most consistently effective solver, strong inductive reasoning, rapid adaptation to emerging evidence, and stable progression across diverse scientific domains. Mid-sized models such as Qwen2.5-7B and Qwen2.5-14B display competitive and robust performance relative to their scale, often exhibiting disciplined evidence-based refinement during exploration. In contrast, larger models (Qwen2.5-72B and DeepSeek-R1-70B) exhibit higher variance: they perform well on specific tasks but frequently show delayed breakthroughs, unstable trajectories, and inconsistent cumulative progress. These patterns suggest that larger models tend to spend more time exploring alternative possibilities rather than leveraging early evidence, resulting in delayed commitment and reduced adaptability during the search process.
3 Discussion
The emergence of the autonomous scientific discovery system (ASDS) represents a paradigm shift in scientific research, with the potential to accelerate discovery by delegating structured experimental processes to artificial intelligence. Yet most existing agentic ML frameworks and benchmark-driven ASDS, such as MLE-bench [99], emphasize an agent’s ability to generate and execute candidate solutions, offering limited support for adaptive trajectory management, such as reasoning about when exploration should continue, how search strategies should evolve, or why certain regions of the space warrant further investigation. SelfAI addresses this gap by integrating structured intent interpretation, multi-agent reasoning, and principled stopping criteria into a general reasoning framework. Across diverse domains, this design enables the system to allocate computational effort where it is most informative, rather than exhaustively enumerating the search space.
Under this goal, SelfAI is meticulously designed as a general-purpose platform capable of performing high-level reasoning across diverse tasks, thereby accelerating the experimental search process. Evaluations across a comprehensive benchmark (12 tasks across 6 different domains) uncover that SelfAI strategically allocates limited computational resources to the most promising performance spaces or research directions and terminates unproductive trajectories, exhibiting human-like cognitive flexibility in dynamically adjusting the search trajectory. This optimal stopping behavior effectively alleviates the most persistent inefficiencies in current ASDS.
The newly proposed Score and $\text{AUP}_{D}$ metrics are crucial for quantifying this advantage, revealing SelfAI’s remarkable balance between rapidly finding discovery efficiency (Score) and maintaining healthy exploration diversity ($\text{AUP}_{D}$). These metrics further illuminate the limitations of traditional methods like grid search and Bayesian optimization, which often lack the adaptive reasoning required to escape local optima or terminate unfruitful trials. Our findings also suggest that in scientific domains, effective scientific discovery demands specialized reasoning capabilities beyond mere model scale. Although large-scale models like GPT-4o and DeepSeek-R1-70B occasionally excel in broad exploration, their performance is frequently surpassed by more compact, purpose-driven models such as Qwen2.5-7B and DeepSeek-R1-7B, which demonstrate superior stopping efficiency and search consistency. This indicates that for navigating complex non-convex optimization environments, the structured reasoning and hypothesis generation implemented in SelfAI’s cognitive agent may play a more decisive role than model scale alone. The framework demonstrates that purpose-driven architectural design can effectively harness the capabilities of moderately-sized models to deliver robust scientific discovery performance.
The implications of this framework extend far beyond accelerated experimental search and improved benchmark performance. First, the architecture of SelfAI embodies a practical blueprint for human-AI symbiosis in science. The User Agent acts as a seamless translator, allowing human researchers to interact with the system at the level of scientific intent and high-level questions, rather than through low-level code or configuration files. This design formally integrates human expertise as a guiding input, creating a collaborative loop where the AI manages the scale and complexity of the search while the human provides strategic direction and domain insight. This addresses a critical shortcoming of "black-box" automation by fostering reproducibility and trust, as every action of the AI is traceable to a human-defined objective. Furthermore, this closed-loop process allows researchers to shift from using AI for scientific research to exploring how to conduct scientific research optimally and to investigate efficient search strategies and problem-related endpoints under controlled conditions.
SelfAI has the potential to evolve from a tool into a collaborative research partner, enhancing the cognitive and experimental capabilities of human scientists. This evolution would be driven by several key advancements. For instance, the User Agent could be enhanced to incorporate autonomous code generation, thereby broadening the range of experimental configurations without requiring explicit user pre-definition. Furthermore, while the current iteration of SelfAI relies on the static knowledge embedded within its underlying LLM (Large Language Model), future versions could leverage advanced memory management and RAG approaches to alleviate context length limitations and dynamically integrate a richer, more up-to-date body of domain-specific knowledge. Collectively, these envisioned capabilities underscore SelfAI’s foundational role in shaping a more adaptive, collaborative, and intelligent paradigm for autonomous scientific discovery.
4 The Design of SelfAI for Scientific Discovery
SelfAI is a general scientific discovery system designed to automate and accelerate the end-to-end scientific discovery process. Its core function is to transform users’ high-level research intentions (hyperparameter optimization, novel algorithm design, model architecture validation, or ablation studies) into structured, executable workflows. The system operates in an integrated loop, comprising three core agents: a user agent, a cognitive agent, and an experiment manager, which together constitute the overall experimental workflow in Fig. 1.
4.1 User Agent
The User Agent, as the user interface of SelfAI, is designed to be interactive, user-controllable, and adaptable to evolving research needs. It interprets high-level scientific intent and translates natural-language objectives into structured, machine-readable experimental configurations. Through iterative clarification, it helps researchers specify objectives, constraints, and search spaces, thereby establishing the experimental environment in which the Cognitive Agent operates. For example, requests such as “achieve state-of-the-art accuracy on CIFAR-10 using a CNN” or “identify the most influential method for image classification” are reformulated into precise experimental specifications (see Supplementary Material A.1). Unlike static intent-generation prompts, the User Agent supports continuous intervention: researchers may pause execution, inspect reasoning traces, adjust constraints, or redesign experimental protocols at any stage without restarting the workflow.
4.2 Cognitive Agent
SelfAI’s cognitive agent is central to accelerating scientific discovery. After receiving structured experimental configurations from the user agent and accumulated experimental history from the experiment manager, the agent synthesizes all available information, analyzes experimental trajectories, evaluates observed performance trends, and identifies promising regions in the search space. This achieves a delicate balance between developing these high-potential regions and exploring areas of greater uncertainty, ultimately deriving the next optimal search strategy. A significant feature of our cognitive agent is its introduction of principled stopping criteria, addressing a key limitation of traditional AI systems, which often continue exploring under conditions of diminishing returns. By conducting evidence-based evaluations of research progress, the agent can determine when further experiments on a particular hypothesis are unlikely to yield significant new insights, thereby reallocating resources to more promising research directions (see Supplementary Material A.2 for more details). This capability improves resource utilization efficiency in the automated discovery process.
4.3 Experiment Manager
The Experiment Manager executes the experimental plans proposed by the Cognitive Agent, each of which has already been structured as executable training jobs, specifying model architectures, datasets, and hyperparameters. For every job, the manager initializes these specified configurations, launches the training and evaluation pipeline, and records all associated outputs. It tracks runtime status, handles execution failures, and captures performance metrics as well as relevant logs and metadata. These records constitute a continuously updated structured knowledge base, providing a reliable foundation for the cognitive agent to analyze search progress, identify promising regions of the search space, and decide when to stop exploration. Comprehensive execution procedures and implementation details are provided in Supplementary Material A.3.
Beyond executing trials, the Experiment Manager also manages the computational environment in which scientific experiments are carried out. Through the Experiment Manager, resources are allocated reliably, execution workloads are scheduled efficiently, and checkpointing and recovery mechanisms are coordinated seamlessly. Anomalies or failed runs are surfaced to the Cognitive Agent as feedback for complex adaptive reasoning. As a result, the overall research process is conducted within an efficient, reliable, and reasoning-driven environment.
Table 1: Comparison of SelfAI with related AI research frameworks and benchmarks across system-level, agent-specific, and task-specific capabilities.
Functions Ours Code LLaMA [53] MLGym [48] AI Co-Scientist [45] AIRA [24] MLAgentBench [49] Optuna [61]
Interactive Research
Flexible Artifacts
Privacy and Security
Trajectory Analysis
Hypothesis Generation
Strategic Planning
Causal Inference
Adaptive Learning
Job Scheduling
Checkpoint Management
Experiment Tracking (*)
Zero-Code Parallelization
Hypothesis Optimization (*)
Self-Evaluation
Benchmark Suite
  • Note: The comparison is structured in four blocks: (1) System-level functions, (2) Cognitive functions primarily handled by the Cognitive Agent, (3) Execution functions managed by the Experiment Manager, and (4) Performance in optimization tasks. (*) denotes basic support.
4.4 Comparison of AI research systems
We contextualize SelfAI’s capabilities within the current landscape of AI-assisted research tools. Table 1 presents a systematic comparison against representative frameworks across multiple dimensions of scientific functionality. Existing ASDSs often specialize in code- or idea-level hypothesis generation under fully autonomous pipelines that primarily optimize medal rates within 24 hours. In contrast, SelfAI integrates intent interpretation, iterative scientific reasoning, and adaptive trajectory optimization into a unified discovery workflow. These capabilities allow the system to refine strategic planning based on accumulated experimental evidence, rather than executing predetermined and static generalization procedures. Overall, these advantages position SelfAI as a general-purpose framework capable of supporting efficient, reliable, and reasoning-driven scientific workflows that extend the scope of existing agent or hyperparameter optimization benchmarks.
5 Data availability
All datasets used in this study are available for download at https://github.com/XiaoXiao-Woo/SelfAI.
6 Code availability
The SelfAI repository is available as Supplementary Software. Updated versions can be found at https://github.com/XiaoXiao-Woo/SelfAI.
Appendix A LLM Agents in SelfAI System for Scientific Discovery
The SelfAI framework supports the complete lifecycle of AI systems, covering design, training, evaluation, and deployment. It integrates three core agents: the User Agent, the Cognitive Agent, and the Experiment Manager, and provides an integrated toolkit for autonomous model training, experiment orchestration, and advanced reasoning, enabling a scalable and adaptive workflow.
A.1 User Agent: Idea Interaction and Experiment Configuration
In autonomous research systems, the User Agent is typically powered by a general-purpose LLM and acts as the primary interface between human researchers and the system. These systems build an “idea-to-experiment” autonomous research process where user access to detailed experimental workflows is limited. Such approaches can streamline research execution but may also restrict researchers’ flexibility in iteratively shaping and adapting experiments. By contrast, our User Agent unifies idea generation and experiment configuration into a continuous, interactive process, empowering researchers to refine initial concepts, explore alternative approaches, and adjust experimental parameters at any stage.
During the idea interaction process, the User Agent understands the task background, dissects method details, and defines precise experimental objectives across diverse scientific domains. At this stage, the User Agent can help users pinpoint innovative research directions and generate field-relevant ideas. Once an idea emerges, the configuration interaction phase leverages prompt templates to clarify user objectives and formulate strategic experimental plans, including initial hypotheses and search configurations. This process converts high-level manual queries into standardized and reproducible experimental configurations, as illustrated in the following examples. We insert the specified template into the prompt and trigger the prompt to fill out the specified configuration file.
Consequently, the User Agent can transform ideas into system and user information. The system information provides the Cognitive Agent with essential content, including role-playing instructions and task-specific knowledge (such as task benchmarking, method evaluation, comparative analysis, ablation studies, and experimental designs). This information equips the Cognitive Agent to interpret high-level research objectives accurately and to execute relevant reasoning and planning steps effectively.
In parallel, the user information focuses on the practical aspects of experimental planning by supplying detailed specifications and keywords, such as search space definitions, the number of trials, optimization criteria, and resource constraints, which are critical for configuring and running experiments. Together, these two types of information enable seamless coordination between conceptual ideation and concrete execution, supporting a workflow in which researchers can iteratively refine their experiments and monitor progress in real time.
User Agent Prompt for Configuration Interaction System: User: {{SUMMARIZED_CONTENT}} Please fill out the content following the YAML format: 1# Template for Configuration Interaction 2# to convert Idea Interaction into YAML format 3- role: system 4 content: 5 model: \$modelName 6 description: You are a \$role specializing in 7 studying \$taskName. Please provide 8 professional and detailed answers. 9 task: \$taskName 10 basic_idea: \$basic_idea 11 search_space: 12 \$hyper_name1: [] 13 \$hyper_name2: [] 14 ... 15 link: \$paperLink 16 instrustion: Complete instructions under limited trials. 17- role: user 18 content: 19 max_trials: \$maxTrials 20 trials: []
A.2 Cognitive Agent
Figure 5: Illustration of the Cognitive Agent. The overall reasoning process involves several key steps: Hypothesis Generation (analysis of the current task and completed trials), Stopping Judgment, and Strategic Planning. Strategic Planning develops experimental schemes based on the analyzed hypotheses.
In our SelfAI framework, the Cognitive Agent continuously ingests and adapts domain knowledge to tackle complex reasoning problems during experiments, much like a human researcher. At each stage, the agent first conducts a comprehensive analysis of the task and completed experiments to clarify the purpose and key information of the task and generates preliminary hypotheses. Then, it evaluates whether these hypotheses are worth trying through stop judgment to inform strategic planning. It then generates specific schemes to probe unexplored regions of the solution space, creating a non-Markovian chain of evolving plans grounded in prior states.
A.2.1 Reasoning for Scientific Discovery
LLM agents exhibit remarkable reasoning capabilities, making them well-suited for scientific discovery. As illustrated in Fig. 5, we present an advanced reasoning framework for generating optimal stopping solutions, an approach that can be extended to various scientific discovery tasks.
A. Trajectory Analysis and Hypothesis Generation
Specifically, the agent begins with Task 1: Analyze the Current Task, where it meticulously interprets the prompt to establish the task’s core objectives, constraints, and hyperparameters. This initial analysis is used to perform a structured trajectory evaluation, creating a coherent chain of thought that guides all subsequent strategy and exploration. This initial analysis is used to perform a structured trajectory evaluation, creating a coherent chain of thought that guides all subsequent reasoning and planning strategies.
Next, in Task 2: Analysis of Completed Trials, the agent systematically evaluates previous outcomes to generate new hypotheses. It specifically analyzes the performance trends of individual parameters and their combinations to identify promising directions in the broader hyperparameter space and propose novel configurations. This deep trend analysis allows the agent to anticipate promising regions of the search space and detect valuable reasoning trajectories that might otherwise be overlooked.
Cognitive Prompt for Trajectory Analysis and Hypothesis Generation System: {{SYSTEM_PART_OF_CONFIGURATION}} User: Completed trials: {{COMPLETED_TRIALS}} Task 1: Analyze the current task Understand current tasks, basic ideas, objectives, and hyperparameters. Task 2: Analysis of Completed Trials Step 1: Summarize performance metrics for completed trials. Step 2: Evaluate performance trends for hyperparameters. Step 3: Highlight promising hyperparameter combinations.
B. Best Stopping Judgement As described in Sect. LABEL:sec:self-evaluation, we introduce the optimal stopping criterion to guide the prompt design that balances exploration and exploitation. During Best Stopping Judgement, the cognitive agent first determines whether the current performance metrics surpass those of the initial configuration and conducts an optimal stopping judgment. Building on the prior analysis of completed trials, the agent systematically evaluates stopping conditions to avoid testing all configurations.
By using prompt instructions, the optimal stopping judgment evaluates all completed trials by leveraging insights from trial analysis, observed performance trends, and identified key findings. In practice, we implement this judgment process via a structured prompt template comprising two main tasks and explicit stopping criteria. The language model then analyzes the relationship between completed trials and unexplored areas. The prompt is organized as follows:
Cognitive Prompt for Best Stopping Judgement User: Completed trials: {{COMPLETED_TRIALS}} The following **Search Space** contains **unexplored** trials. {{TRIALS}} Instructions: Task 1: Review Analysis of Completed Trials (trial analysis, performance trends, highlights, and other insights) Task 2: Decide Whether to Stop Optimization Based on the above analysis and **Completed Trials**, determine whether the optimization process should be stopped. Carefully analyze each of the following stop rules and provide a short (1-2 sentences) justification for whether it is met: 1. Have all promising configurations identified based on performance trends been tested? 2. Is it unlikely that unexplored configurations will perform better based on the observed trends and the law of diminishing returns? 3. Has the best metric improved significantly? Step 2: Decide whether **all** conditions are met. If **all** criteria in Step 1 are met, Answer: Yes, with confidence score: {{CONFIDENCE_SOCRE}}. Otherwise, Answer: No with confidence score: {{CONFIDENCE_SOCRE}}.
C. Strategic Planning Following the reasoning and stopping judgment phases shown in Fig. 5, the agent reviews all executed trials to synthesize key observations on performance trends, notable configurations, and areas of consistent improvement. Next, it translates its abstract hypotheses into a concrete experimental plan. This strategic process balances two competing objectives: exploiting promising configurations and exploring uncertain regions. Exploitation focuses on refining or building upon previously successful configurations, whereas exploration targets under-examined areas that show significant potential for performance gains. Each proposed trial is explicitly justified by insights from the preceding reasoning phase, ensuring the search is deliberate and well-founded. Finally, each recommendation is formalized as an intended trial and delegated to the Experiment Manager for execution.
As trials advance and incremental improvements diminish, the Experiment Manager maintains a dynamic record of completed actions and continuously refines the remaining search space. Strategic planning directs the optimization trajectory toward convergence, and the system employs an effective stopping strategy to terminate exploration in a rational and efficient manner.
Notably, the LLM-based agent demonstrates human-like exploratory behavior: it initially focuses on refining hypotheses around high-performing regions and gradually extends attention to areas of greater uncertainty. This pattern reflects a synergy of intelligence-driven reasoning and adaptive planning.
Cognitive Prompt for Strategic Planning User: Instructions: Task 1: Review Analysis of All Completed Trials Completed trials: {{COMPLETED_TRIALS}} The following **Search Space** contains **unexplored** trials: {{TRIALS}} Instructions: Task 2: Optimization Recommendation Recommend exactly {{N_JOBS}} promising trials from the provided **Search Space** (include both number and params). Rules: 1. “params” MUST include: {{HYPERNAME}} 2. All selected ‘params‘ must match exactly with the provided **Search Space**. Do NOT leave out any key. 3. Use the analysis in **Task 1** (trial analysis, performance trends, highlights, and other insights) to guide selection. 4. Based on the above analysis, explore under-explored regions only when there is clear evidence of potential performance gain. 5. Do not mix, modify, or create new values. 6. You MUST not output any JSON blocks in this part. 7. You MUST provide reasoning for each recommendation.
A.3 Experiment Manager
The Experiment Manager is responsible for experiment orchestration and recovery, including resource management, task allocation, and progress tracking. These capabilities enable efficient coordination of multi-instance parallel optimization, maximize resource utilization, and enhance training robustness:
1) Resource Management. The Experiment Manager monitors user program resource consumption and dynamically allocates available GPU, TPU, and memory resources. This granular allocation optimizes workload distribution across computing units and ensures stable execution of all trials.
2) Fault recovery and Checkpoint Reconnection. In case of system interruptions or suboptimal model performance, the Experiment Manager reports failures to the cognitive agent. The Experiment Manager performs preliminary diagnostics, identifies potential issues, adjusts training parameters, and resumes training from the latest checkpoint.
3) Multi-Instance Parallel Optimization. SelfAI instantiates each user program to run across diverse physical environments, independent of the target program framework. The Experiment Manager coordinates multi-instance parallel training, synchronizes execution, and concurrently tests various configurations, thereby shortening overall training time and improving generalization across datasets and model parameters. For each parallel experiment, the Experiment Manager identifies and supplies necessary runtime parameters, ensuring experiments are conducted under the same environment optimized by the cognitive agent.
SelfAI maintains detailed logs of all configurations, training runs, and evaluation metrics, and provides targeted optimization suggestions for each completed experiment. This intelligent decision-making refines experimental design and reduces redundant search in automated scientific discovery. The collaboration between the Experiment Manager and the cognitive agent establishes a self-improving workflow. By conducting advanced reasoning and analyzing trial result patterns through hypothesis-driven exploration, they uncover relationships between settings and model performance, steering the search toward optimal configurations. This ensures that parallel training is not merely a brute-force process, but an intelligent exploration of the training space that adapts based on real-time outcomes.
A.4 Evaluation for Reasoning Trajectories
While the reasoning process of LLMs in problem-solving often generates a diverse range of discrete insights and multiple potential chains of thought, this diversity, while valuable for exploration and exploitation, can pose a challenge to coherent reasoning evaluation across various experimentation and discovery phases. We propose a systematic evaluation metric that captures both the diversity of reasoning perspectives and the overall coherence of the reasoning process, ensuring a more robust and comprehensive assessment of reasoning capabilities.
Optimal Stopping Criteria In this work, we collected the best value point and the stop point from trials. Based on Optimal Stopping Criteria [100, 101], we can define the following measure formulas,
$\text{Gain}=\frac{1}{N}\sum^{N-1}_{i=0}\frac{v^{*}_{i}-v_{i,\min}}{v_{i,\max}-v_{i,\min}}$ (1)
$t_{\text{best}}=\frac{1}{N}\sum^{N-1}_{i=0}t^{\text{best}}_{i}=\frac{1}{N}\sum^{N-1}_{i=0}\frac{m_{i}}{M_{i}}$ (2)
$t_{\text{stop}}=\frac{1}{N}\sum^{N-1}_{i=0}t^{\text{stop}}_{i}=\frac{1}{N}\sum^{N-1}_{i=0}\frac{n_{i}}{M_{i}}$ (3)
and where $N$ is the number of tasks. For the $i$-th task, $M_{i}$ is the number of completed trials. $m_{i}$ is the best value point index. $t_{i}^{\text{best}}$ is the cost of the best value. In addition, we set the stop point index, $n_{i}$, where $t_{i}^{\text{stop}}$ is better when $n_{i}$ is closer to the best value point index. $S_{i}$ means that the binarized value.
To obtain a comprehensive measure, we combine the last three measures, i.e., Rel, $t_{\text{best}}$, and $t_{\text{stop}}$, where we utilize Rel in underestimated penalty, then $t_{\text{best}}$ and $t_{\text{stop}}$ are the time penalty ($P_{\text{best}}$ and $P_{\text{stop}}$). Thus, the total penalty is
$P_{\text{total}}$ $=\frac{t_{i}^{\text{stop}}+t_{i}^{\text{best}}}{2}$ (4)
Finally, the score is denoted as
$\text{Score}=\frac{1}{N}\sum^{N-1}_{i=0}\text{Gain}\cdot(1-P_{total})$ (5)
Best Approximation/Candidate In [48], performance profiles and the AUP aim to measure available rates across m tasks, where all performance metrics are threshold $\tau$, performance profiles ($\rho_{\texttt{m}}(\tau)$) are computed as thresholds in all metrics (sorted by ascending) in current task.
$\text{AUP}_{\texttt{m}}=\int^{\tau_{\max}}_{1}\rho_{\texttt{m}}(\tau)d\tau$ (6)
It is noted that the above performance profile and $\text{AUP}_{\texttt{m}}$ score cannot measure the diversity of reasoning. Therefore, we rewrite the performance profile and AUP score:
First, the performance profile is defined in all completed trials $M_{i}$ in $i$-th task and the overall search space $\mathcal{H}$, as follows
$r_{i}=\begin{cases}\frac{\max{\{v_{k}:\,k\in\mathcal{H}\}}}{v_{i}},\,\text{ascend}\\ \\ \frac{v_{i}}{\min{\{v_{k}:\,k\in\mathcal{H}\}}},\,\text{descend}\end{cases}$ (7)
where ascend/descend denotes that the value is larger/smaller, the performance is better. For all trials, $\tau$ is the set of all obtained $r_{i}$ values.
Then, we consider all completed trials $M_{i}$ in $i$-th task,
$\rho_{i}(\tau)=\begin{cases}|\{k\in M_{i}:r_{k}>=\tau\}|,&\text{ascend}\\ |\{k\in M_{i}:r_{k}<=\tau\}|,&\text{descend}\end{cases}$ (8)
which captures how many evaluated configurations exceed a given performance threshold $\tau$. $\rho_{i}(\tau)$ is the cumulative distribution curve of the trajectory.
The area term aggregates the overall concentration of strong configurations along the performance axis:
$A=\frac{1}{N}\sum^{N-1}_{i=0}\int_{\tau_{\min}}^{\tau_{\max}}\rho_{i}(\tau)d\tau.$ (9)
To capture the temporal asymmetry of discovery, we compute the centroid
$G=\frac{1}{A}\int_{\tau_{\min}}^{\tau_{\max}}x\cdot\rho_{i}(\tau)\,d\tau,$ (10)
and define the skewness
$S=\int_{\tau_{\min}}^{\tau_{\max}}\left(\frac{x}{G}\right)^{3}\rho(x)\,d\tau.$ (11)
Since $S$ may be unbounded and may take both positive (left-skewed) and negative (right-skewed) values, we normalize it via a reference skewness value $S_{\text{base}}$ from the GS method and a smooth monotonic mapping:
$S’=1-\frac{S-S_{\text{base}}}{S_{\text{base}}},$ (12)
$S^{\prime}=\frac{\tanh(S)+1}{2},S^{\prime}\in(0,1).$ (13)
Finally, the Area Under the Performance–Diversity curve ($\text{AUP}_{D}$) is defined as
$\text{AUP}_{D}=A/S^{\prime},$ (14)
where trajectories that exhibit earlier concentration of high-performing configurations obtain larger and thus smaller $\text{AUP}_{D}$ value, whereas trajectories that concentrate improvements later yield smaller and therefore larger value.
Appendix B Details of SelfAI Benchmark
This section provides the complete technical definitions of the metrics used in the main text, including all normalization procedures, ranking aggregation, and scoring rules. These details are omitted from the main text for clarity and conciseness.
Table 2 summarizes the 12 tasks spanning six scientific categories used to evaluate the reasoning capacity of different solvers. For each dataset, we supply background information to the LLMs to simulate the user-intent interpretation process (see Sect. A.1). Most datasets are generated through systematic grid-based exploration, whereas LCBench and the Chagas EP20 drug-discovery dataset are constructed from Bayesian optimization trajectories. LCBench, in particular, is a widely used AutoML benchmark  [71]. Supplementary Fig. 6 provides an additional structural analysis of the LCBench search space. Accuracy depends on interacting hyperparameters such as weight decay, hidden-layer width, and depth, and the high-performance region is sharply localized. This structure highlights why solvers must possess strong reasoning capabilities to efficiently locate these promising configurations. Supplementary Fig. 7 reports solver rankings based on the Score metric. These rank summaries allow cross-task comparisons of solver consistency; detailed interpretation of performance trends is provided in the main text.
In Table 3, we summarize averaged performance across all benchmark tasks. Classical baselines such as grid search and TPE-based Bayesian optimization achieve competitive final objective values but receive low Score and high $\text{AUP}_{D}$, reflecting redundant late-stage exploration and delayed stopping. Naive LLM solvers improve early discovery but remain unstable across tasks. In contrast, SelfAI-driven solvers based on mid-sized models such as Qwen2.5-7b and GPT4-o3-mini attain the highest Scores with substantially lower $\text{AUP}_{D}$ and earlier stopping, indicating more focused and efficient trajectories. Larger models (e.g., Qwen2.5-72b, Llama3.3-70b) do not necessarily yield better efficiency, highlighting that structured reasoning and trajectory control within SelfAI are more critical than sheer model scale. In addition, as discussed in Sect. B.1, failure analyses are also included for completeness. These examples illustrate typical breakdowns in reasoning processes across long trajectories, such as premature stopping, limited incorporation of earlier observations, or sensitivity to minor perturbations. They inform the broader evaluation of solver stability and support the claims presented in the main text.
Table 2: List of tasks in SelfAI with different hyperparameters for multiple tasks and datasets.
Category Method Task Dim Count Ref.
Scientific Computing Tensor Network Image Completion 3 64 [65]
Random Forest Regression 5 162 [62]
Deep Learning LSTM Sentiment Analysis 2 20 [63]
GraphSAGE Node Classification 22 25 [93]
Computer Vision SIREN Image Denoising 2 25 [68]
SIREN Image Segmentation 2 25 [68]
ResNet Image Classification 4 9 [70]
MAE Image Classification 2 20 [69]
FashionMnist-NN Image Classification 5 2000 [71]
Medical Image Analysis nnUnet BraTS [82] 3 18 [72]
nnUnet-revisited BTCV [81] 5 19 [73]
Drug Discovery DNN Bioactivity Prediction 4 30 [94]
Figure 6: The relationship between different parameter settings and accuracy. a A hexbin plot showing the joint density of learning rate (log-scaled) and accuracy. b Box plots illustrating the distribution of dropout rates used for models of different depths. c Mean accuracy and standard error across logarithmically binned batch sizes. d Performance comparison of different architectural configurations (layers and units). e A scatter plot of weight decay against momentum, colored by accuracy. f The optimization trajectory, showing accuracy improvement over successive trials, with a moving average trend line. g A correlation matrix quantifying linear relationships between all parameters and the accuracy.
Table 3: Averaged performance comparison of SelfAI across different tasks.
[…]
GS 0.2453 1.0000 0.5094 1.0000 1.0000 1.0000 14
BS 0.1927 0.8106 0.6265 0.9881 1.0000 1.0000 15
LLM 0.3526 0.7638 0.2949 1.0000 1.0000 0.9286 13
LLM-ES 0.5294 0.2349 0.4691 0.4582 0.9981 0.6429 3
Qwen2.5-7b 0.5562 0.2154 0.4805 0.3945 0.9957 0.7857 2
Qwen2.5-14b 0.5015 0.3310 0.4926 0.4997 0.9969 0.7857 5
Qwen2.5-32b 0.4287 0.4684 0.5252 0.6133 0.9972 0.7857 10
Qwen2.5-72b 0.4189 0.5358 0.4531 0.7087 0.9995 0.8571 11
DeepSeek-r1-7b 0.4769 0.1802 0.7020 0.3093 0.9927 0.5000 8
DeepSeek-r1-14b 0.4793 0.3956 0.4953 0.5100 0.9948 0.7143 7
DeepSeek-r1-32b 0.4989 0.3476 0.4535 0.5433 0.9933 0.7857 6
DeepSeek-r1-70b 0.4556 0.3513 0.5392 0.5299 0.9962 0.7143 9
Llama3.3-70b 0.3625 0.5483 0.5099 0.7271 0.9683 0.7143 12
GPT4-o3-mini 0.6433 0.2259 0.3168 0.3961 0.9989 0.8571 1
GPT4-o3 0.5140 0.2284 0.5477 0.3961 0.9966 0.6429 4
Figure 7: Performance ranking of different methods across multiple tasks. The heatmap displays the rank of each method (rows) for every task (columns), where lower numbers (darker colors) indicate better performance (e.g., 1st place). The final average rank is summarized on the right.
BS
Refer to caption Refer to caption Refer to caption Refer to caption
GPT4-o3-mini
Refer to caption Refer to caption Refer to caption Refer to caption
GPT4-o3
Refer to caption Refer to caption Refer to caption Refer to caption
Llama3.3-70B
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 8: Illustration of the optimized trajectory using the SIREN method for additional cases in image denoising (first two columns) and segmentation (last two columns). Blue points are initial points. Green points represent suggested points before reaching the optimum. Red points indicate redundant suggestions generated after the optimal point has been reached. The $\star$ marks the optimal point. The numbered labels indicate the sequence of recommendations provided by LLMs.
Qwen2.5-7B
Refer to caption Refer to caption Refer to caption Refer to caption
Qwen2.5-14B
Refer to caption Refer to caption Refer to caption Refer to caption
Qwen2.5-72B
Refer to caption Refer to caption Refer to caption Refer to caption
Qwen2.5-72B
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 9: More optimized trajectories of Fig. 8. These LLMs have different suggestions for non-convex hyperparameter optimization.
DeepSeek-R1-7B
Refer to caption Refer to caption Refer to caption Refer to caption
DeepSeek-R1-14B
Refer to caption Refer to caption Refer to caption Refer to caption
DeepSeek-R1-32B
Refer to caption Refer to caption Refer to caption Refer to caption
DeepSeek-R1-70B
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 10: More optimized trajectories of Fig. 8. These LLMs have different suggestions for non-convex hyperparameter optimization.
B.1 Failure Cases
Comparison of Best Result Hit Rate. Table 3 compares the experimental results of traditional solvers and LLM-based solvers. Early-stopping LLM-ES partially alleviates redundant trials but remains less stable and less efficient than SelfAI-integrated variants. In contrast, SelfAI consistently enhances trajectory-level reasoning across all LLM families, reducing unnecessary exploration (lower $\text{AUP}_{D}$), improving stopping behavior (lower Stop-Time), and yielding higher overall Score. Mid-sized models such as Qwen2.5-7B and DeepSeek-R1-32B demonstrate particularly favorable exploration and exploitation trade-offs. Although LLMs possess advanced reasoning capabilities and can identify optimal results more efficiently, we found that LLM-based solvers still face challenges related to context limitations and computational fragility.
1) Context limitations. The LCBench dataset exceeds the maximum token limit of GPT-family models, which is why SelfAI does not employ GPT4-o3-mini or GPT4-o3. In practice, search spaces rarely contain thousands of trials, and this issue can be mitigated in future work. For instance, memory management and Retrieval-Augmented Generation (RAG) can help reduce token requirements by introducing additional steps to the experimental recommendation workflow, such as embedding, chunking, and retrieval.
2) Computational fragility. Table 3 compares the best result hit rates across solvers. The performance of LLM-based solvers is highly sensitive to the reasoning strategy, and early stopping often leads to a decline in hit rate. In the supplementary tables, we report the best result $t_{\text{best}}$ for each solver. $t_{\text{best}}=1$ indicates that the optimal result was not found. Moreover, when $t_{\text{best}}=1$, the score metric improves significantly, highlighting the solver’s weakness in optimal stopping. As illustrated in Fig. 11, computational fragility also hinders reliable performance improvement through parameter tuning, as seen in the non-monotonic performance of the DeepSeek-R1 model series. Table 6 summarizes the metrics corresponding to Fig. 11.
Fig. 11 provides a detailed visualization of solver behavior. The Qwen2.5 models consistently identify high-performing configurations early in the search, with Qwen2.5-7B reaching near-optimal regions within the first 10–15 trials. Their trajectories exhibit stable refinement and limited oscillation, suggesting consistent integration of prior experimental evidence. In contrast, the DeepSeek-R1 models show pronounced non-monotonic behavior. DeepSeek-R1-7B frequently oscillates between high and low values, reflecting sensitivity to minor variations in its reasoning process. DeepSeek-R1-70B achieves early improvements but then terminates prematurely, failing to further refine promising regions. Although the GS and BS reach high-performance regions, their trajectories lack the rapid breakthroughs observed in LLM-based solvers. Overall, these trajectory comparisons provide visual evidence of differences in exploration stability, breakthrough efficiency, and stopping behavior across solver families, complementing the aggregate metrics reported in the main text.
Figure 11: Failure Cases for the DeepSeek-R1 Family in the scientific computing field. Both DeepSeek-R1-7B and DeepSeek-R1-70B fail to reach the optimal results, marked with a star symbol ($\star$). Notably, DeepSeek-R1-7B fails to improve beyond the initial results, as indicated by the red dotted line, suggesting limited exploration capability in the search space.
Appendix C Supplementary Tables
Table 4: Performance comparison of different solvers for the Boston housing prices based on the Random Forest Regressor.
[…]
GS 0.4969 1.0000 0.0062 1.0000 0.841 9
BS 0.4654 0.9745 0.0755 0.9937 0.841 11
Llama3.3-70b 0.3831 0.1330 1.0000 0.1509 0.837 13
Qwen2.5-7b 0.7893 0.3463 0.0943 0.3270 0.841 5
Qwen2.5-14b 0.7013 0.6108 0.0063 0.5912 0.841 7
Qwen2.5-32b 0.6981 0.4081 0.2579 0.3459 0.841 8
Qwen2.5-72b 0.7862 0.5250 0.0126 0.4151 0.841 6
DeepSeek-R1-7b 0.9057 0.0910 0.0881 0.1006 0.841 3
DeepSeek-R1-14b 0.8522 0.3401 0.0063 0.2893 0.841 4
DeepSeek-R1-32b 0.9403 0.1618 0.0063 0.1132 0.841 2
DeepSeek-R1-70b 0.3948 0.0215 1.0000 0.0189 0.833 12
GPT4-o3-mini 0.9811 0.0062 0.0189 0.0189 0.841 1
GPT4-o3 0.4694 0.0560 1.0000 0.0377 0.840 10
Table 5: Performance comparison of different solvers for the sentiment analysis task based on the LSTM model.
[…]
GS 0.4750 1.0000 0.0500 1.0000 0.96 2
BS 0.0294 0.8817 0.9412 1.0000 0.96 13
Llama3.3-70b 0.3824 0.8817 0.2353 1.0000 0.96 4
Qwen2.5-7b 0.1830 0.2857 1.0000 0.5294 0.94 12
Qwen2.5-14b 0.3529 0.8318 0.4118 0.8824 0.96 7
Qwen2.5-32b 0.3529 0.6383 0.5882 0.7059 0.96 8
Qwen2.5-72b 0.4412 0.8220 0.2353 0.8824 0.96 3
DeepSeek-R1-7b 0.2157 0.1419 1.0000 0.3529 0.93 11
DeepSeek-R1-14b 0.2353 0.8817 0.5294 1.0000 0.96 10
DeepSeek-R1-32b 0.3824 0.8817 0.2353 1.0000 0.96 4
DeepSeek-R1-70b 0.3824 0.8817 0.2353 1.0000 0.96 4
GPT4-o3-mini 0.8824 0.1494 0.0588 0.1765 0.96 1
GPT4-o3 0.2745 0.0469 1.0000 0.1765 0.93 9
Table 6: Performance comparison of different solvers for scientific computing fields, where we evaluate the tensor wheel decomposition method on the multi-spectral image completion dataset.
[…]
GS 0.2656 1.0000 0.4688 1.0000 41.50 11
BS 0.1066 0.9638 0.7869 1.0000 41.50 13
Llama3.3-70b 0.4016 0.8979 0.2623 0.9344 41.50 8
Qwen2.5-7b 0.7377 0.2904 0.2295 0.2951 41.50 1
Qwen2.5-14b 0.5410 0.2779 0.4262 0.4918 41.50 2
Qwen2.5-32b 0.2869 0.9638 0.4262 1.0000 41.50 10
Qwen2.5-72b 0.1803 0.9638 0.6393 1.0000 41.50 12
DeepSeek-R1-7b 0.3216 0.3009 1.0000 0.3443 41.21 9
DeepSeek-R1-14b 0.4426 0.8687 0.2295 0.8852 41.50 4
DeepSeek-R1-32b 0.4426 0.6645 0.4426 0.6721 41.50 4
DeepSeek-R1-70b 0.4136 0.1477 1.0000 0.1639 41.34 7
GPT4-o3-mini 0.5410 0.2623 0.4590 0.4590 41.50 3
GPT4-o3 0.4217 0.0539 1.0000 0.1475 41.34 6
Table 7: Performance comparison of different solvers for the image segmentation task based on the SIREN model.
[…]
GS 0.1100 1.000000 0.780000 1.000000 16.63 13
BS 0.329545 0.761200 0.477273 0.863636 16.63 11
Qwen2.5-7b 0.602273 0.487619 0.363636 0.431818 16.63 2
Qwen2.5-14b 0.454545 0.708497 0.272727 0.818182 16.63 4
Qwen2.5-32b 0.374388 0.538240 0.613636 0.636364 16.60 9
Qwen2.5-72b 0.408757 0.536178 0.636364 0.545455 16.60 5
DeepSeek-r1-7b 0.693182 0.318656 0.227273 0.386364 16.63 1
DeepSeek-r1-14b 0.408534 0.471791 0.704545 0.477273 16.60 6
DeepSeek-r1-32b 0.312477 0.198417 1.000000 0.340909 15.77 12
DeepSeek-r1-70b 0.374555 0.576330 0.590909 0.659091 16.60 8
Llama3.3-70b 0.380344 0.285865 0.590909 0.295455 10.79 7
GPT4-o3-mini 0.477273 0.517709 0.454545 0.590909 16.63 3
GPT4-o3 0.363636 0.671366 0.477273 0.795455 16.63 10
Table 8: Performance comparison of different solvers for the image denoising task based on the SIREN model.
[…]
GS 0.1500 1.0000 0.7000 1.0000 24.78 12
BS 0.1477 0.8819 0.7045 1.0000 24.78 13
Qwen2.5-7b 0.7045 0.1802 0.1364 0.4545 24.78 4
Llama3.3-70b 0.5706 0.1105 0.5455 0.2955 24.27 8
Qwen2.5-14b 0.4205 0.5205 0.5455 0.6136 24.78 10
Qwen2.5-32b 0.5795 0.4762 0.3182 0.5227 24.78 7
Qwen2.5-72b 0.7614 0.1543 0.1364 0.3409 24.78 1
DeepSeek-R1-7b 0.5341 0.4193 0.3864 0.5455 24.78 9
DeepSeek-R1-14b 0.7159 0.1639 0.0909 0.4773 24.78 2
DeepSeek-R1-32b 0.7159 0.1284 0.2045 0.3636 24.78 2
DeepSeek-R1-70b 0.6818 0.1359 0.2955 0.3409 24.78 5
GPT4-o3-mini 0.2703 0.4389 0.7955 0.6591 24.48 11
GPT4-o3 0.6023 0.4786 0.2045 0.5909 24.78 6
Table 9: Performance comparison of different solvers for Mask AutoEncoders (MAE) [69].
[…]
GS 0.1000 1.0000 0.8000 1.0000 85.0 12
BS 0.0294 0.8093 0.9412 1.0000 85.0 13
Llama3.3-70b 0.2941 0.8093 0.4118 1.0000 85.0 10
Qwen2.5-7b 0.7647 0.0112 0.2353 0.2353 85.0 2
Qwen2.5-14b 0.3471 0.1513 1.0000 0.2941 84.5 9
Qwen2.5-32b 0.2647 0.8093 0.4706 1.0000 85.0 11
Qwen2.5-72b 0.4412 0.8093 0.1176 1.0000 85.0 7
DeepSeek-R1-7b 0.4104 0.0080 1.0000 0.1765 84.9 8
DeepSeek-R1-14b 0.6176 0.3853 0.2941 0.4706 85.0 5
DeepSeek-R1-32b 0.6765 0.1108 0.2941 0.3529 85.0 3
DeepSeek-R1-70b 0.5588 0.2235 0.1765 0.7059 85.0 6
GPT4-o3-mini 0.8235 0.0088 0.1176 0.2353 85.0 1
GPT4-o3 0.6471 0.1141 0.3529 0.3529 85.0 4
Table 10: Performance comparison of different solvers for the image classification task on the ImageNet dataset reported from the ResNet benchmark.
[…]
GS 0.0000 1.0000 1.0000 1.0000 21.43
BS 0.0833 0.6394 0.8333 1.0000 21.43
Llama3.3-70b 0.3333 0.6394 0.3333 1.0000 21.43
Qwen2.5-7b 0.6667 0.0865 0.3333 0.3333 21.43
Qwen2.5-14b 0.8333 0.0000 0.1667 0.1667 21.43
Qwen2.5-32b 0.5833 0.2192 0.3333 0.5000 21.43
Qwen2.5-72b 0.4167 0.2988 0.5000 0.6667 21.43
DeepSeek-R1-7b 0.4167 0.3322 0.5000 0.6667 21.43
DeepSeek-R1-14b 0.6667 0.0865 0.3333 0.3333 21.43
DeepSeek-R1-32b 0.7500 0.0865 0.1667 0.3333 21.43
DeepSeek-R1-70b 0.5000 0.2443 0.5000 0.5000 21.43
GPT4-o3-mini 0.5000 0.2192 0.5000 0.5000 21.43
GPT4-o3 0.5833 0.2192 0.3333 0.5000 21.43
Table 11: Performance comparison of different solvers for the image classification task on the LCBench dataset reported from [71].
[…]
GS 0.1750 1.0000 0.6500 1.0000 88.29
BS 0.4189 0.9985 0.1622 1.0000 88.29
Llama3.3-70b 0.4738 0.0052 1.0000 0.0486 87.98
Qwen2.5-7b 0.4619 0.0047 1.0000 0.0451 85.78
Qwen2.5-14b 0.4739 0.0017 1.0000 0.0180 85.62
Qwen2.5-32b 0.4718 0.0022 1.0000 0.0225 85.62
Qwen2.5-72b 0.4599 0.0081 1.0000 0.0766 87.98
DeepSeek-R1-7b 0.4691 0.0027 1.0000 0.0280 85.62
DeepSeek-R1-14b 0.4579 0.0086 1.0000 0.0806 87.98
DeepSeek-R1-32b 0.4677 0.0035 1.0000 0.0331 85.78
DeepSeek-R1-70b 0.4739 0.0014 1.0000 0.0180 85.62
GPT4-o3-mini - - - - - -
GPT4-o3 - - - - - -
Table 12: Performance comparison of different solvers for the medical image segmentation task on the BTCV dataset reported from the nnUnet benchmark [73].
[…]
GS 0.4211 1.0000 0.1579 1.0000 85.04 3
BS 0.0625 0.6453 0.8750 1.0000 85.04 13
Llama3.3-70b 0.3750 0.6453 0.2500 1.0000 85.04 5
Qwen2.5-7b 0.4688 0.3441 0.5000 0.5625 85.04 2
Qwen2.5-14b 0.2813 0.6453 0.4375 1.0000 85.04 7
Qwen2.5-32b 0.2813 0.4742 0.6875 0.7500 85.04 7
Qwen2.5-72b 0.1875 0.6453 0.6250 1.0000 85.04 11
DeepSeek-R1-7b 0.3680 0.0813 1.0000 0.1250 83.69 6
DeepSeek-R1-14b 0.1983 0.0505 1.0000 0.1875 80.69 10
DeepSeek-R1-32b 0.1875 0.6453 0.6250 1.0000 85.04 11
DeepSeek-R1-70b 0.2813 0.6453 0.4375 1.0000 85.04 7
GPT4-o3-mini 0.6875 0.1466 0.2500 0.3750 85.04 1
GPT4-o3 0.4063 0.5081 0.4375 0.7500 85.04 4
Table 13: Performance comparison of different solvers for the nnUnet model on the BraTS dataset [82].
[…]
GS 0.1667 1.0000 0.6667 1.0000 82.45 10
BS 0.0333 0.3786 0.9333 1.0000 82.45 13
Llama3.3-70b 0.1333 0.3786 0.7333 1.0000 82.45 11
Qwen2.5-7b 0.1000 0.3786 0.8000 1.0000 82.45 12
Qwen2.5-14b 0.4333 0.0698 0.5333 0.6000 82.45 1
Qwen2.5-32b 0.3333 0.3118 0.5333 0.8000 82.45 3
Qwen2.5-72b 0.2333 0.0883 0.7333 0.8000 82.45 9
DeepSeek-R1-7b 0.2995 0.0481 1.0000 0.4000 82.41 5
DeepSeek-R1-14b 0.2940 0.0406 1.0000 0.4000 82.00 7
DeepSeek-R1-32b 0.2667 0.0839 0.6667 0.8000 82.45 8
DeepSeek-R1-70b 0.3333 0.3126 0.5333 0.8000 82.45 3
GPT4-o3-mini 0.4000 0.2347 0.6000 0.6000 82.45 2
GPT4-o3 0.2984 0.0484 1.0000 0.4000 82.33 6
Table 14: Performance comparison of different solvers for the GraphSAGE model on the imbalanced node classification task.
[…]
GS 0.1000 1.0000 0.8000 1.0000 89.28 13
BS 0.2727 0.8982 0.4545 1.0000 89.28 8
Llama3.3-70b 0.1591 0.8982 0.6818 1.0000 89.28 9
Qwen2.5-7b 0.2880 0.0928 1.0000 0.4091 89.12 6
Qwen2.5-14b 0.2894 0.0890 1.0000 0.4091 89.15 5
Qwen2.5-32b 0.2880 0.3564 1.0000 0.4091 89.12 6
Qwen2.5-72b 0.1364 0.8982 0.7273 1.0000 89.28 12
DeepSeek-R1-7b 0.3116 0.3083 1.0000 0.3636 89.15 3
DeepSeek-R1-14b 0.1591 0.8982 0.6818 1.0000 89.28 9
DeepSeek-R1-32b 0.1591 0.8657 0.7273 0.9545 89.28 11
DeepSeek-R1-70b 0.2955 0.7096 0.5909 0.8182 89.28 4
GPT4-o3-mini 0.7727 0.1463 0.0455 0.4091 89.282 1
GPT4-o3 0.7727 0.2540 0.1818 0.2727 89.282 1
Table 15: Performance comparison of different solvers for bioactivity prediction on Chagas EP20 dataset [95]
[…]
GS 0.4833 1.0000 0.0333 1.0000 0.754 11
BS 0.3333 0.8941 0.3333 1.0000 0.754 13
Llama3.3-70b 0.4630 0.8941 0.0741 1.0000 0.754 12
Qwen2.5-7b 0.9074 0.0772 0.0741 0.1111 0.754 1
Qwen2.5-14b 0.8889 0.0649 0.1111 0.1111 0.754 2
Qwen2.5-32b 0.6296 0.4231 0.0741 0.6667 0.754 9
Qwen2.5-72b 0.5741 0.6799 0.0741 0.7778 0.754 10
DeepSeek-R1-7b 0.7778 0.1104 0.2222 0.2222 0.754 4
DeepSeek-R1-14b 0.7037 0.5510 0.0741 0.5185 0.754 7
DeepSeek-R1-32b 0.6852 0.3405 0.0741 0.5556 0.754 8
DeepSeek-R1-70b 0.7778 0.3160 0.1111 0.3333 0.754 4
GPT4-o3-mini 0.7407 0.3546 0.1852 0.3333 0.754 6
GPT4-o3 0.8148 0.0622 0.0370 0.3333 0.754 3
Appendix D Prompts of Compared Methods
Search Prompt of LLM solver Instructions: **Search Space** (Numbering starts from 0, excluding Completed Trials): {{TRIALS}} Task 1: Optimization Recommendation Recommend exactly {{N_JOBS}} promising trials from the provided **Search Space** (include both number and params). Rules: 1. “params” MUST include: {{HYPERNAME}} 2. All selected ‘params‘ must match exactly with the provided **Search Space**. Do NOT leave out any key. 3. Use the analysis in **Task 1** (trial analysis, performance trends, highlights, and other insights) to guide selection. 4. Based on the above analysis, explore under-explored regions only when there is clear evidence of potential performance gain. 5. Do not mix, modify, or create new values. 6. You MUST not output any JSON blocks in this part. 7. You MUST provide reasoning for each recommendation.
Early-Stopping Prompt of LLM-ES solver Completed trials: {{COMPLETED_TRIALS}} The following **Search Space** contains **unexplored** trials. {{TRIALS}} If the optimization process should be stopped, Answer: Yes with confidence score: {{CONFIDENCE_SOCRE}}. Otherwise, Answer: No with confidence score: {{CONFIDENCE_SOCRE}}. Finally, you MUST output ’Answer: No/Yes’ with confidence score: {{CONFIDENCE_SOCRE}}.

### v2

@@
-SelfAI: Building a Self-Training AI System with LLM Agents
+SelfAI: A self-directed framework for long-horizon scientific discovery
 Abstract
-Recent work on autonomous scientific discovery has leveraged LLM-based agents to integrate problem specification, experiment planning, and execution into end-to-end systems.
-However, these frameworks are often confined to narrow application domains, offer limited real-time interaction with researchers, and lack principled mechanisms for determining when to halt exploration, resulting in inefficiencies, reproducibility challenges, and under-utilized human expertise.
-To address these gaps, we propose SelfAI, a general multi-agent platform that combines a User Agent for translating high-level research objectives into standardized experimental configurations, a Cognitive Agent powered by LLMs with optimal stopping criteria to iteratively refine hyperparameter searches, and an Experiment Manager responsible for orchestrating parallel, fault-tolerant training workflows across heterogeneous hardware while maintaining a structured knowledge base for continuous feedback.
-We further introduce two novel evaluation metrics, Score and $\text{AUP}_{D}$, to quantify discovery efficiency and search diversity.
-Across regression, NLP, computer vision, scientific computing, medical imaging, and drug discovery benchmarks, SelfAI consistently achieves strong performance and reduces redundant trials compared to classical Bayesian optimization and LLM-based baselines, while enabling seamless interaction with human researchers.
-1 Main
-In recent years, large language models (LLMs) [1, 2, 3] have fundamentally reshaped the landscape of AI-driven research.
-Advances in reasoning [4, 5], multimodal understanding [6, 7], and autonomous tool use [8, 9, 10, 11, 12, 13] have positioned LLMs as central components for accelerating scientific workflows.
-Early LLM-based scientific research demonstrated that LLMs can extract actionable scientific knowledge to guide experiments [14, 15, 16, 17] or answer professional questions [18, 19, 20].
-As reasoning and tool-interaction capabilities matured, research expanded to cross-stage planning [21, 22, 23], hypothesis and idea generation [24, 25, 26, 27], multimodal knowledge integration [28], and conducting experiments [17, 29].
-Meanwhile, LLMs have driven the development of tasks such as reaction prediction [30, 31, 32], material discovery [33, 34, 18, 35], chemical synthesis design and molecular property optimization [36, 29, 37, 30], and biomedicine research [17, 38, 39, 40, 41].
-Building on these advances, a new generation of scientific discovery systems has emerged, spanning AI scientific assistants across the full spectrum of scientific workflows [18, 42, 16, 43], automated scientific discovery systems (ASDSs) specialized for code generation and experiment design [36, 44, 29], and increasingly capable research agents that incorporate automated debugging, iterative idea refinement, and scientific writing [26, 45, 46].
-Along this evolutionary trajectory, LLM-centered ASDS have begun to integrate inference, execution, and feedback mechanisms into unified closed-loop pipelines, achieving varying degrees of autonomy across scientific domains.
-Despite rapid progress, existing ASDS frameworks still exhibit fundamental limitations.
-Many systems [44, 45, 24, 47] (such as AIRA and Scientist-V2) excel at translating research intent into executable code or experimental procedures, thereby validating the “executability” of ASDS and primarily relying on final success rate (medal rate) or best performance as metrics.
-More autonomous platforms like MLGym [48] and MLAgentBench [49] enable system-level control, allowing agents to execute end-to-end benchmarks, collect results, and conduct standardized evaluations.
-Moreover, while recent LLM-driven optimization methods [50, 51, 52, 53] such as LLM4EO showcase that LLMs can infer evolutionary tendencies and synthesize more effective operators, they operate at the level of local modification rules rather than scientific reasoning in trajectories.
-Overall, these approaches do not explicitly enhance scientific reasoning during experimentation, such as identifying optimal stopping points, discovering efficient trial trajectories, or assessing the structural quality of reasoning.
-Building reasoning-centric ASDS therefore requires a shift from execution-oriented autonomy toward cognitive autonomy: systems must not only carry out experiments but also analyze, reflect, and revise their reasoning throughout the experimental process.
-However, current ASDS systems provide little support for these cognitive-layer behaviors.
-The absence of mechanisms for trajectory-level reasoning assessment and LLM-driven decision-making limits both efficiency and accuracy, especially in settings with large search spaces, stochastic outcomes, or complex interdependencies among hypotheses, resulting in inefficient exploration, elevated computational costs, and limited real-time adaptability.
-These limitations highlight a fundamental need for ASDS architectures that incorporate reasoning-driven search, adaptive learning mechanisms, and structural evaluation of scientific trajectories.
+Scientific discovery increasingly entails long-horizon exploration of complex hypothesis spaces, yet most existing approaches emphasize final performance while offering limited insight into how scientific exploration unfolds over time, particularly balancing efficiency-diversity trade-offs and supporting reproducible, human-in-the-loop discovery workflows.
+We introduce SelfAI, a self-directed, multi-agent-enabled discovery system that automates scientific exploration as a strategic, trajectory-driven decision-making process.
+SelfAI translates high-level research intent into executable experiments, reasons over accumulated experimental trajectories to guide subsequent exploration, and applies adaptive stopping decisions to terminate unproductive search paths within a closed-loop workflow governed by explicit efficiency-diversity trade-offs.
+Evaluated using real-world experiments spanning domains from machine learning to drug discovery, SelfAI consistently discovers high-quality solutions with substantially fewer redundant trials than classical optimization and recent LLM-based baselines.
+The proposed methods establish a general framework for organizing long-horizon scientific discovery and adaptive decision-making in complex scientific and engineering systems.
+1 Introduction
+Scientific discovery across engineering [bradshaw1983studying, angelopoulos2024transforming], physics [lupoiu2025multi, kaiser2025large], biology [lyu2024alphafold2, nguyen2024sequence], and medicine [yakavets2025machine, grisoni2021combining, jiang2022artificial] increasingly relies on artificial intelligence (AI) systems to support or partially automate key stages of the scientific workflow, ranging from data analysis and hypothesis generation to experimental design and execution.
+As scientific problems grow in scale, dimensionality, and uncertainty, computational methods have become indispensable for exploring complex hypothesis spaces.
+AI-assisted discovery systems hold the promise of fundamentally changing the way scientific discovery is practiced [xin2025towards].
+Most existing AI-assisted discovery systems [aira, yakavets2025machine, erps2021accelerated] adopt paradigms rooted in machine learning, emphasizing predictive modeling on large datasets and evaluating success primarily through final performance metrics.
+While effective for many applications, these paradigms provide limited support for reasoning about the discovery process itself, including how hypotheses are explored, how experiments are prioritized, and how decisions are made under uncertainty.
+In particular, they rarely treat scientific discovery as a sequential, long-horizon decision process in which exploration strategies, resource allocation, and stopping decisions must be jointly optimized over time.
+Unlocking the full potential of AI in discovery, therefore, requires moving beyond performance-centric automation toward frameworks that explicitly support exploration, decision-making, and adaptation throughout the discovery trajectory.
+Recent advances in large language models (LLMs) [openai2023gpt, bai2023qwen, guo2025deepseek] have significantly expanded the potential of AI systems in scientific research.
+Improvements in reasoning, multimodal understanding, and autonomous tool use now enable LLM-based systems to integrate planning, execution, and feedback within unified workflows.
+As a result, these systems can support key components of scientific discovery, including hypothesis generation, experiment design, and iterative refinement across domains [huang2022towards, ferrag2025llm].
+Building on these capabilities, prior work has progressively enhanced individual aspects of the scientific workflow.
+Early efforts demonstrated that LLMs can extract actionable scientific knowledge to guide experiments [lin2023evolutionary, angelopoulos2024transforming] and answer domain-specific professional questions [alampara2025probing, polak2024extracting, dagdelen2024structured].
+Subsequent research extended these capabilities to cross-stage planning [biomni], automated hypothesis generation [wang2025discovery, aira, baek2025researchagent], multimodal knowledge integration [steyaert2023multimodal, gao2025chemical], and automated experimentation across scientific domains.
+These advances have led to a new generation of scientific discovery systems, ranging from AI research assistants to fully autonomous workflows capable of executing end-to-end benchmarks [king2004functional, mandal2025evaluating, ai_co_scientist].
+While these systems demonstrate the feasibility of system-level automation in scientific workflows, most existing approaches focus primarily on improving individual stages of the discovery process or optimizing final outcomes.
+Even recent LLM-driven optimization methods [liao2025llm4eo, jiang2025agenticsciml, kochnev2025optuna], which infer evolutionary patterns and synthesize update rules, operate largely at the level of execution or local improvement.
+However, scientific discovery is inherently a sequential and strategic process that unfolds over time.
+It requires reasoning not only about isolated steps, but about the structure and quality of the entire exploration trajectory.
+Current systems rarely treat discovery as a trajectory-level decision-making problem.
+Consequently, they do not explicitly address fundamental challenges such as determining optimal stopping points, designing efficient exploration paths, or evaluating the structural effectiveness of discovery strategies over time.
+Addressing these limitations requires a shift from execution-oriented autonomy toward cognitively oriented autonomy, in which discovery systems reason about exploration strategies, evaluate efficiency-diversity trade-offs, and adaptively decide when to stop.
+Motivated by this perspective, we introduce SelfAI, a self-directed, multi-agent-enabled discovery system that treats scientific discovery as a trajectory-level decision-making process.
+We integrate human intent, exploration reasoning, and experimental execution into a unified closed-loop framework.
+These are built with three modules and external tools (Fig.
+1a), enabling efficiency-diversity exploration in long-horizon scientific workflows.
+To evaluate discovery quality, we introduce two novel complementary metrics, Score (Discovery Efficiency) and $\text{AUP}_{D}$ (Area Under the Performance Diversity).
+Score aggregates, across tasks, the normalized improvement over the search space together with penalties for discovering good configurations late and for stopping far from the best-found point.
+$\text{AUP}_{D}$ explicitly encodes how broadly a solver explores the search space by summarizing the performance-diversity tradeoffs across the entire trajectory, enabling detailed analysis of exploration behavior and stopping decisions in long-term searches.
+Together, these metrics measure quantitative assessments of exploration behavior, reasoning structure, and stopping decisions in long-term autonomous experiments.
 Figure 1: SelfAI Framework for Automated Scientific Experimentation.
 a, Holistic architecture of the multi-agent system, which transforms various experiments in the research process into a structured workflow.
-b, User intentions, comprising ideas and experiment schemes, are transformed into structured configurations via a predefined prompt.
+b, User intents, comprising ideas and experiment schemes, are transformed into structured configurations via a predefined prompt.
 These inputs are processed through successive stages: hypothesis generation, strategic planning, trial execution, and result collection.
-c, Performance distribution across 11 tasks demonstrates the framework’s ability, when powered by GPT-4o-mini, to prioritize high-performance regions without sacrificing exploration.
-d, Trial counts for each solver shown in c, accompanied by quantile lines, density distributions, and performance variability across the global and two evaluation regions.
-Higher values in low-performance regions promote rapid escape, while lower values in high-performance regions enable localized refinement.
-To address these issues, we propose SelfAI, a unified multi-agent self-training pipeline for continuous adaptability and transparent collaboration.
+As illustrated in Fig.
+1b, SelfAI operates as a closed-loop discovery system that unifies research intent, trajectory-aware reasoning, and a diverse set of executive experiment tools to support various scientific problems and long-horizon scientific discovery.
+The workflow begins with a User Agent, which translates high-level research objectives and exploratory questions into standardized experiment configurations.
+These structured configurations provide a consistent interface across heterogeneous environments, enabling reliable comparison and reuse of results.
+Cognitive Agent evaluates trial outcomes, reasons over accumulated discovery trajectories, and adaptively revises exploration strategies over time.
+Experimental execution is coordinated by Experiment Manager, which autonomously manages resource allocation, environment provisioning, checkpoint-based execution, and comprehensive experiment logging.
+Through continual feedback between reasoning and execution, the Cognitive Agent and Experiment Manager jointly assess the expected utility of continued exploration and apply adaptive stopping decisions to terminate unproductive search paths.
+In this way, SelfAI enables sustained improvement through long-horizon, trajectory-driven discovery.
+In summary, SelfAI provides a self-directed framework that treats scientific discovery as a strategic, trajectory-driven process.
+By iteratively integrating intent specification, trajectory-aware reasoning, and adaptive stopping, SelfAI enables principled allocation of experimental resources while preserving exploration diversity under realistic computational and experimental constraints.
+Across a broad range of scientific discovery tasks, SelfAI consistently achieves efficient, diverse, and high-quality exploration in heterogeneous and large-scale search spaces.
+These results demonstrate that trajectory-aware reasoning and adaptive stopping allow discovery systems to reduce redundant experimentation, identify diminishing returns early, and redirect effort toward more informative regions of the search space.
+More broadly, this work highlights the role of LLM-guided decision-making in organizing scientific exploration under explicit efficiency-diversity trade-offs.
+Figure 2: Performance comparison and efficiency-diversity trade-offs across solvers.
+a, Normalized performance of five solvers (GS, BS, LLM, LLM-ES, and SelfAI) evaluated by Score (higher is better), $\text{AUP}_{D}$, Best-time, and Stop-time (lower is better).
+Bars indicate solver-level averages, with values in parentheses denoting the best-achieved or optimal reference performance.
+b, Outcome distributions across solvers.
+Left: kernel density estimates illustrating score concentration and convergence tendencies.
+Right: violin plots with overlaid individual trials (sample size n shown), capturing performance variability.
+The dashed horizontal line denotes the evaluation threshold separating low- and high-performance regimes; annotated statistics (red/blue) highlight solver behavior above and below this threshold, where broader exploration facilitates rapid escape from low-performance regions and tighter distributions in high-performance regions enable localized refinement.
+2 Results
+2.1 SelfAI: Long-horizon scientific discovery
+SelfAI is a trajectory-aware paradigm for long-horizon scientific discovery (Fig.
+1).
+It treats accumulated experimental outcomes as structured evidence within an evolving discovery trajectory, from which exploration decisions, hypothesis refinement, and stopping judgments are iteratively and autonomously derived.
+SelfAI builds on recent LLM-based optimization and agentic frameworks that interleave reasoning with trial proposals but typically emphasize short-horizon improvement or final performance.
+Across experiments, SelfAI consistently synthesizes accumulated information, analyzes experimental trajectories, evaluates performance trends, and prioritizes promising regions of the search space.
+2.2 Evaluation of long-horizon scientific discovery
+To characterize efficiency-diversity trade-offs in long-horizon exploration, we propose four metrics: Score, $\text{AUP}_{D}$, $t_{\text{best}}$, and $t_{\text{stop}}$ (Fig.
+2a), the details of which are described in Sect.
+4.4.
+These metrics not only reflect final performance but also explicitly capture when high-quality solutions are discovered, how broadly the search space is explored, and how efficiently exploration is terminated.
+Score aggregates normalized improvement while penalizing both delayed discovery of optimal configurations and prolonged exploration after performance plateaus.
+$\text{AUP}_{D}$ (Area Under the Performance-Diversity Curve) quantifies the diversity of explored high-quality solutions and summarizes how performance gains are distributed along the discovery trajectory.
+The metrics $t_{\text{best}}$ and $t_{\text{stop}}$ denote the normalized time first to identify the best result and the time at which exploration terminates, respectively.
+An effective solver minimizes both quantities, achieving early discovery while avoiding unnecessary trials.
+In this framework, effective exploration strategies attain a high Score with a relatively low $\text{AUP}_{D}$, indicating efficient discovery with minimal wasted exploration.
+Detailed per-task comparisons of Score, $\text{AUP}_{D}$, $t_{\text{best}}$, and $t_{\text{stop}}$ are provided in the Supplementary Materials B.
+2.3 SelfAI reorganizes exploration dynamics across scientific domains
+We evaluate long-horizon scientific discovery using a benchmark constructed from real-world scientific discovery tasks (Materials and Methods) and assess solver performance under these efficiency-diversity trade-offs (Fig.
+2).
+Using the same underlying language model (GPT-4o) across all solvers, we observe that while most methods are eventually able to identify high-performing configurations, their exploration dynamics differ substantially.
 As shown in Fig.
-1, SelfAI forms a closed loop that integrates user intent, cognitive reasoning, and experimental orchestration.
-Upstream, a User Agent converts high-level objectives and exploratory questions into standardized experiment configurations.
-These configurations enable the Cognitive Agent to analyze performance metrics, conduct reasoning over historical outcomes, and adjust the search trajectory, driving gradual improvement [54, 55].
-A resource-aware Experiment Manager ensures efficient execution by handling resource scheduling, environment provisioning, adaptive checkpointing, and comprehensive experiment logging.
-During high-level parameter optimization, the Cognitive Agent and Experiment Manager jointly estimate the potential benefit of continued training and apply an optimal-stopping criterion to terminate unpromising trials.
-To evaluate scientific reasoning quality, we introduce two complementary evaluation metrics, Score (Discovery Efficiency) and $\text{AUP}_{D}$ (Area Under the Performance Diversity).
-Score aggregates, across tasks, the normalized improvement over the search space together with penalties for discovering good configurations late and for stopping far from the best-found point.
-$\text{AUP}_{D}$ explicitly encodes how broadly a solver explores the search space by summarizing the performance-diversity tradeoffs across the entire trajectory, enabling detailed analysis of exploration behavior and stopping decisions in long-term searches.
-Together, these metrics enable quantitative assessment of exploration behavior, reasoning structure, and stopping decisions in long-term autonomous experiments.
-In summary, SelfAI provides a continuously self-training system that autonomously iterates through cycles of intent-centric understanding, trajectory-aware reasoning, and adaptive strategic planning, enabling sustained performance gains with minimal human intervention.
-Through checkpoint tolerance, zero-code parallelization, and dynamic resource management, SelfAI substantially reduces redundant computation, improves resource efficiency, and accelerates search convergence in long-horizon scientific workflows.
-In addition, SelfAI integrates seamlessly with modern MLOps pipelines [56, 57, 58], facilitating collaborative experiment coordination, end-to-end traceability of model evolution, and robust privacy protection for sensitive configurations and user data.
-Collectively, these capabilities make SelfAI a practical, scalable, and operationally mature platform for rapid scientific discovery.
-The key contributions of this paper are given as follows:
-  1. 1.
-    Autonomous Scientific Discovery System: We design a cohesive framework of specialized LLM agents, including the User Agent, Cognitive Agent Framework, and Experiment Manager, which collaboratively translate high-level research objectives into iteratively refined experimental workflows, including trajectory analysis, hypothesis generation, and best stopping decision-making.
-  2. 2.
-    Novel Evaluation Metrics and Optimal Stopping: We propose two new metrics, Score (discovery efficiency) and $\text{AUP}_{D}$ (area under the performance–diversity curve), to jointly quantify the efficiency and diversity of hyperparameter exploration.
-An embedded optimal stopping criterion automatically terminates unpromising trials, effectively balancing exploration and exploitation while minimizing computational overhead.
-  3. 3.
-    Comprehensive Validation of LLM-based Reasoning in Scientific Discovery: We evaluate SelfAI across diverse tasks, including regression, sentiment analysis, computer vision, medical imaging, and drug discovery.
-The results show that SelfAI effectively improves upon excellent baseline models in both results exploration and performance.
-2 Performance of SelfAI on Benchmark Data
-Figure 2: Scores among all solvers across different tasks to measure the best stopping criterion.
-Figure 3: Diverse Metrics ($\text{AUP}_{D}$) among all solvers across different tasks to evaluate trajectory diversity.
-Scientific discovery is driven by diverse physical environments and user intents.
-To effectively reason across diverse scientific discovery tasks, our approach integrates multiple LLM agents and a suite of tools: User Agent, Cognitive Framework, and Experiment Management, addressing the complex reasoning and planning problem of scientific discovery.
-In our experiments, we evaluate the SelfAI across 6 domains and 12 tasks.
-The compared methods contain different suites of LLMs, such as OpenAI-o3 [1], Llama3.3 [59], Qwen2.5 [2], and DeepSeek-R1 [3].
-We used grid search and the Tree Parzen Estimator (TPE) optimizer [60, 61] (referred to as BS), which applies a modified Bayesian inference method using a spiral search.
-All agents are tested on four Nvidia A100 GPUs.
-A public implementation of SelfAI will be available on the project website.
-We also provide reasoning processes for reproducibility.
-[…]
-Benchmarks.
-We collect benchmarks composed of multiple tasks of 6 primary categories, designed to simulate challenging scenarios in scientific discovery.
-The benchmark requires the solver to engage in scientific reasoning by analyzing experiments, interpreting observations, and adapting strategies, uncovering novel patterns, functional relationships, or optimal experimental configurations.
-Compared with hyperparameter search, these processes reflect the inductive reasoning essential to real-world scientific inquiry.
-Therefore, this benchmark assesses the model’s reasoning ability, not just its numerical optimization capabilities.
-All benchmark data is collected through actual experimental runs, and some results are reported in published literature, spanning both discrete and continuous hyperparameter search spaces, from low-dimensional to high-dimensional settings (see Appendix B for more details).
-Existing methods [24, 45, 49] explore final performance or medal rates using reward functions and prompts specifically designed for innovative idea generation, code generation, and pass@K strategies.
-In the line of LLM-driven optimization, LLM4EO [50] uses LLMs exclusively to synthesize and evolve evolutionary operators that determine which parts of a candidate solution should be modified during the search.
-The LLM is used to generate and update operators to shape the local search behavior of the evolutionary algorithm.
-These implementations do not focus on scientific reasoning across trials during the exploration process, making direct comparisons with our benchmark infeasible.
-Consequently, we used the readily accessible GPT4-o3-mini model and compared SelfAI with LLM and LLM-ES methods on 11 small/medium-sized tasks across different domains in Fig.
-1c.
-The LLM and LLM-ES methods are fine-tuned by using the structured prompts in Supplementary Section D.
-The LLM solver uses a search prompt that recommends new trials based on prior results, while LLM-ES augments this with an early-stopping module that decides whether to terminate exploration.
+2a, SelfAI consistently outperforms LLM and LLM-ES baselines on $\text{AUP}_{D}$, Best-Time, and Stop-Time, indicating reduced redundant experimentation and more effective allocation of trials toward informative regions of the search space.
+Distributional analyses (Fig.
+2b) further reveal that SelfAI reallocates experimental effort away from diminishing-return regions earlier while maintaining sufficient coverage of uncertain regions.
+Across the benchmark, for low-performance regimes, SelfAI obtains higher variances, conducts fewer trials to escape low-performance regimes, and concentrates effort in promising regions with lower variances.
+Finally, SelfAI achieves highly efficient exploration with 168 runs by terminating exploration near the best-achieved configuration under fixed trial budgets.
+Benchmark construction details and additional analyses are provided in Fig.
+5.
+2.3.1 Case studies
+We present representative case studies to illustrate how trajectory-aware reasoning and adaptive stopping manifest in concrete scientific discovery scenarios, illustrating how different solvers organize exploration trajectories over time, allocate experimental effort, and determine when further experimentation becomes uninformative.
+In addition, we illustrate that trajectory-level metrics capture discovery efficiency, diversity, and how reasoning limitations can emerge in structured optimization settings.
+Additional task-specific analyses, including 12 tasks, are provided in Supplementary Sect.
+B.
+Machine learning.
+We first evaluate SelfAI on the Boston housing price prediction task [boston] using a random forest regression model.
+As shown in Fig.
+3 and Supplementary Table 4, nearly all solvers achieve a similar and high best result ($\approx 0.841$), indicating that differences in final predictive accuracy are minimal.
+Small- and medium-sized models outperform other solvers, but while large models DeepSeek-R1-70B and Llama3.3-70B demonstrated consistent and sustained exploration capabilities, their stopping behavior was less efficient, leading to longer search trajectories and delayed convergence.
+These results illustrate the importance of trajectory-aware reasoning and adaptive stopping in practical scientific optimization, as larger-scale models may exhibit either overly diffuse exploration or premature stopping and are not always aligned with task-specific objectives.
+Further examples are presented below in this section.
+Computer vision.
 Fig.
-1c presents that our SelfAI outperforms LLM and LLM-ES in all four metrics, validating SelfAI’s effectiveness in integrating domain knowledge to accelerate scientific discovery.
-Furthermore, we focus on the reasoning capability of the original LLM model because it is difficult to be fair when comparing different methods.
+4 illustrates these trajectory-level dynamics in the hyperparameter search landscape for image segmentation using SIREN (see also Supplementary Figs.
+7-9).
+The smoothed response surface reveals several obvious peaks, indicating multiple competing high-performance regions.
+Starting from identical initial points, the tree-based Bayesian optimizer follows a spiral-like trajectory that underutilizes promising regions and exhibits limited global coverage, frequently converging to suboptimal local minima.
+In contrast, LLM-based solvers infer trends from accumulated evaluations and adapt exploration accordingly, resulting in broader coverage of the search space and more reliable identification of globally competitive configurations.
+Importantly, these solvers also monitor trajectory-level progress and terminate exploration once improvements plateau.
+This behavior directly translates into higher scores and lower $\text{AUP}_{D}$, demonstrating how trajectory-aware metrics capture both discovery efficiency and controlled diversity beyond final performance alone.
+Medical image analysis.
+We introduce medical image segmentation using nnU-Net [nnunet], exploring performance optimization and best practices in medical image segmentation across 19 representative CNN-, Transformer-, and Mamba-based models.
+These results provide controlled settings for evaluating long-horizon optimization behavior.
+As shown in Fig.
+3 and Supplementary Tables 12-13, trajectory-aware metrics reveal substantial differences in solver behavior that are not apparent from final segmentation accuracy alone.
+In particular, on the BraTS [brats] and BTCV [btcv] datasets, small- and medium-sized models achieve higher scores and lower $\text{AUP}_{D}$, reflecting their ability to rapidly identify high-performing configurations and terminate exploration.
+Trajectory-aware metrics concentrate exploration earlier in high-potential regions and avoid prolonged exploration once performance gains saturate.
+Conversely, solvers with comparable final performance but lower scores exhibit broader and more redundant exploration trajectories, leading to delayed stopping and reduced discovery efficiency.
+Scientific computing.
+Tensor decomposition [hackbusch2014numerical, verstraete2008matrix] is a fundamental tool in scientific computing for high-dimensional data analysis, with broad applications in data compression [li2025lossless], computational acceleration [berezutskii2025tensor], and multimodal data fusion [steyaert2023multimodal].
+The associated optimization problem [tw] presents a structured landscape characterized by competing objectives between compression quality and computational efficiency.
+Effective solvers must therefore balance refinement of promising configurations with controlled exploration of alternative regimes.
+As shown in Supplementary Table 6, GPT4-o3-mini performed well and ranked highly.
+However, Qwen2.5-72b may have over-relied on general reasoning ability and failed to balance mathematical knowledge and reasoning ability.
+Among the DeepSeek series models, although the DeepSeek-R1-7b model failed to find the optimal solution, its search strategy was acceptable, and its score was moderate.
+Although the DeepSeek-R1-70b and GPT-4o models have excellent search diversity and early stopping capabilities, their scores dropped.
+These results reveal that some large-scale models exhibit either overly diffuse exploration or premature stopping for task-specific objectives.
+This case illustrates that expressive reasoning capacity alone is insufficient for effective scientific discovery.
+Instead, successful optimization requires trajectory-level alignment between exploration strategy, stopping decisions, and problem structure.
+Supplementary analyses further document representative success and failure patterns across tasks (Supplementary Sect.
+B.6).
+In particular, hit rates for identifying globally optimal configurations vary substantially across models, with premature stopping limiting convergence in some cases.
+Moreover, increased model scale does not systematically improve long-horizon inference: larger models often sustain extended hypothesis exploration, leading to delayed commitment and greater performance variability under fixed experimental budgets.
+In contrast, several medium-sized models exhibit more stable trajectory organization, reinforcing the role of disciplined inference and adaptive stopping in efficient scientific discovery.
+Result summary Across 12 heterogeneous tasks and language-model backbones (Supplementary Sect.
+B), these results demonstrate that SelfAI yields consistent trajectory-level benefits.
+SelfAI does not merely improve convergence accuracy but fundamentally reorganizes exploration dynamics in long-horizon scientific discovery.
+By reallocating experimental effort away from diminishing-return regions toward structurally informative areas of the search space, SelfAI implements a policy-adaptive exploration strategy that aligns with identifiable principles of long-horizon scientific discovery.
+2.4 Emergent principles of long-horizon scientific discovery
+We further investigate whether long-horizon scientific discovery exhibits consistent, task-agnostic principles when guided by trajectory-aware reasoning.
 Fig.
-1d shows that SelfAI promptly focuses on exploiting higher-performance regions while maintaining a balanced distribution for effective exploration.
-Figs.
-2 and 3 summarize the benchmark performance based on these two metrics.
-The benchmarks show that SelfAI consistently achieves excellent results across various domains and tasks for different LLM models.
-Detailed comparisons covering Score, $\text{AUP}_{D}$, Best-Time, Stop-Time, and Hit Rate are provided in Supplementary Table 3.
-Failure cases.
-Several key failure modes of LLM-based solvers are shown in Supplementary Section B.1.
-First, the hit rate for identifying the best result varies substantially across models, with some solvers (particularly DeepSeek-R1 series) failing to reach the global optimum due to premature stopping.
-Second, limited context windows cause incomplete reasoning over long trajectories, preventing solvers from fully incorporating earlier experimental signals.
-Third, computational fragility leads to non-monotonic performance under seemingly minor perturbations in reasoning strategy.
-These findings indicate that while SelfAI improves overall efficiency, reasoning-driven exploration remains sensitive to solver stability and stopping behavior.
-2.1 Machine Learning
-Boston house pricing prediction.
-We evaluate SelfAI on the Boston housing price prediction task [62] using a random forest regression model.
-A total of 162 trials were conducted across a search space defined by five hyperparameters: i.e., $\text{n-estimators}=[100,200,300]$, $\text{max-depth}=[\text{None},10,20]$, $\text{min-samples-split}=[2,5,10]$, min-samples-leaf=[1,2,5] and $\text{max-features}=[\text{``sqrt'',``log2''}]$).
-As shown in Figs.
-2 and 3 (and Supplementary Table 4), GPT-4o-mini achieves by far the highest Score (0.9811) and the 1st ranking, indicating outstanding optimization efficiency.
-Within the DeepSeek-R1 family, the 32B and 7B variants secure 2nd and 3rd places respectively, demonstrating the best balance between rapid convergence (low Best-Time) and reasonable search diversity among open-source models.
-In contrast, DeepSeek-R1-70B and Llama3.3-70B exhibited consistent exploration but lacked effective stopping criteria, leading to longer convergence times.
-Notably, nearly all models achieved an identical and high Best Result ( 0.841), indicating that while most can find a correct solution, they differ substantially in optimization efficiency and reliability.
-Sentiment analysis.
-We perform experiments for sentiment analysis that focus on identifying opinions, emotions, and attitudes expressed in text.
-All experiments are found in [63] with standard experimental settings, where the pre-trained Word2Vec embeddings [64] are used as input features.
-An LSTM network is then employed to model sentence sequences, converting them into a multi-class classification problem.
-This setup serves to evaluate the agent’s reasoning and optimization capabilities in textual domains.
-As shown in Supplementary Table 5, GPT-4o-mini outperforms all other LLMs, achieving the highest Score (0.8824) by discovering the optimal configuration extremely early while maintaining reasonable search diversity.
-DeepSeek-R1 series and Qwen2.5 series exhibit significantly lower Scores, with many failing to converge quickly or occasionally missing the global optimum.
-The largest frontier model GPT-4o also underperforms markedly (Score 0.2745, rank 11), suggesting that scale alone is insufficient for effective iterative hyperparameter reasoning in recurrent neural architectures.
-2.2 Scientific Computing for Image Completion
-For scientific computing fields, we selected the tensor decomposition method [65].
-The method is an important tool for high-dimensional data analysis and is crucial in applications such as data compression [66], computational acceleration [67], and multi-modal data fusion [28].
-All LLMs are provided with identical mathematical knowledge about TW decomposition generated by GPT-4o.
-As shown in Supplementary Table 6, Qwen2.5 (7b and 14b) and GPT4-o3-mini perform well and rank highly.
-Qwen2.5-72b, on the other hand, may overly rely on general reasoning capabilities and fail to balance mathematical knowledge and reasoning capabilities.
-In the DeepSeek series, although the DeepSeek-R1-7b model can’t find the optimal solution, its search strategy was acceptable with a moderate Score, demonstrating its ability to perform mathematical reasoning.
-DeepSeek-R1-70b and GPT-4o, despite excellent search diversity and early stopping, received a mediocre score, reflecting that their exploration strategy was not well aligned with the optimization goal of tensor decomposition, possibly failing to find the optimal performance between data compression and computational efficiency.
-2.3 Computer Vision
-A.
-SIREN We employed SIREN (Sinusoidal Representation Networks) [68] to evaluate our framework on image segmentation and denoising tasks.
-Leveraging sine activations and coordinate-based inputs, SIREN excels at representing high-frequency signals through continuous implicit representations, making it widely used in scientific computing and physics-based problems.
-However, its performance is highly sensitive to hyperparameters (e.g., learning rate and regularization strength, etc.), requiring careful tuning per dataset.
-The unsupervised nature of SIREN further increases the risk of training instability or divergence with improper settings, making it a strong test case for SelfAI’s hyperparameter optimization capability.
-Fig.
-4 illustrates hyperparameter search trajectories for image segmentation using SIREN (see also Supplementary Figs.
-8, 9 and 10), where the surface, smoothed from the original steep, multi-peak data, reveals distinct search behaviors.
-Given the same three initial points, the tree-structured Bayesian optimizer (BS) follows a spiral trajectory that underutilizes promising starting regions and fails to explore broadly, often converging to local minima.
-In contrast, LLM-based optimizers infer trends from evaluated points, incorporate causal reasoning, and explore more broadly to efficiently locate the global optimum.
-They also monitor progress and halt when improvements plateau, a capability absent in traditional methods.
-Quantitative results (Supplementary Tables 78) reveal distinct task-dependent performance profiles.
-In segmentation, DeepSeek-R1-7B achieves the highest Score (0.693) and ranks first, followed by Qwen2.5-7B and GPT-4-o3-mini.
-Larger models such as Qwen2.5-72B and Llama-3.3-70B rank only mid-range, indicating that scale alone does not ensure effective optimization.
-In denoising, the landscape shifts: Qwen2.5-72B delivers the best performance (Score 0.761), while DeepSeek-R1-14B and DeepSeek-R1-32B jointly occupy second place.
-Notably, models that perform well in segmentation tasks do not necessarily perform well in denoising tasks, suggesting that the effectiveness of the solver largely depends on the degree of matching between its inference strategy and the task.
-B.
-Image Classification We also benchmark two commonly used supervised learning methods in computer vision: Masked Autoencoder (MAE) [69] and ResNet [70].
-These represent fundamentally different learning paradigms where hyperparameter optimization plays a crucial role in achieving the latest performance.
-Mask Autoencoder (MAE) MAE introduces two key hyperparameters: training strategies (linear detection vs.
-fine-tuning) and masking rate, which are explored over a range: $[0.10,0.90]$ with an interval of 0.1.
-The masking rate directly affects the difficulty of the reconstruction task and the quality of the learning representation, while the training strategy determines how to adapt to the downstream task.
-In our experiments (see Supplementary Table 9 for more details), GPT-4o-mini achieved explicit performance with efficiency, demonstrating the extraordinary ability to identify the best mask configuration and training strategies.
-Qwen2.5-7B achieves early convergence and disciplined stopping.
-The DeepSeek-R1 series shows consistent scaling behavior: performance improves from 7B to 14B to 32B, but slightly drops at 70B.
-In traditional solvers, GS and BS lag significantly behind LLM-based solvers, highlighting the advantages of intelligent hyperparameter optimization methods.
-Apart from Qwen2.5-14b and DeepSeek-R1-7b, most solvers can find the optimal solution, but significant differences exist among the various methods in terms of search efficiency and convergence speed.
-ResNet Family On the ImageNet ResNet hyperparameter benchmark, LLMs in the 7B–32B range demonstrate superior optimization efficacy.
-The top-performing solvers, such as Qwen2.5-14B (1st), DeepSeek-R1-32B (2nd), and Qwen2.5-7B with DeepSeek-R1-14B (tied 3rd), consistently identify the optimal ResNet configuration within one or two trials, halting immediately with near-perfect sample efficiency.
-In contrast, larger variants from the same model families drop to middle or lower rankings, consuming over half the search budget with notably higher exploration overhead.
-While GPT-series models, as large-scale counterparts, still achieve competitive results, the 7B–32B class exhibits a clear advantage in balancing accuracy and efficiency.
-For the architectural search on ImageNet, we explored key dimensions such as depth and bottleneck design.
-Qwen2.5-14B attains theoretical peak performance, with its 7B and 32B versions completing further exploration efficiently.
-These results reflect an ability to identify ideal network configurations with minimal wasted exploration.
-LCBench We evaluate SelfAI on the standard AutoML benchmark [71] using Bayesian optimization over 2000 rounds of hyperparameter search, covering critical parameters including learning rate, batch size, network depth, and dropout rate.
-As shown in Supplementary Table 11, DeepSeek-R1-70B, Qwen2.5-14B, and Llama3.3-70B secure the top three positions in the ranking, with DeepSeek-R1-70B attaining the highest overall score and the lowest search diversity.
-Notably, leading models in the 7B–32B parameter range efficiently identify near-optimal hyperparameter configurations by evaluating only a minimal set of candidate solutions, achieving an order-of-magnitude improvement in stop-time compared to classical methods.
-These findings demonstrate diminishing returns with scaling, as larger models within the same series (e.g., Qwen2.5-72B) typically underperform comparable models of medium size, while the 7B model can achieve robust performance comparable to models several times its size.
+4 illustrates these trajectory-level dynamics in the hyperparameter search landscape for image segmentation using SIREN (Sect.
+2.3.1).
+From principles of long-horizon scientific discovery, trajectory-level dynamics follow a characteristic three-stage trajectory: rapid escape from low-efficiency regions, early concentration of effort in high-potential areas, and termination near the optimal configuration.
+This structured progression highlights the central role of strategic trial allocation and adaptive stopping in the efficiency of long-horizon exploration.
+Aggregated performance across tasks and domains reveals additional trajectory-level regularities (Fig.
+3).
+Despite task-dependent variations in relative rankings (Supplementary Fig.
+6), SelfAI consistently achieves high Scores while maintaining relatively low $\text{AUP}_{D}$ values across heterogeneous tasks and language-model backbones.
+These results indicate that efficient discovery trajectories can be organized according to shared principles that are not tied to any single task or domain.
+Model scale alone does not systematically improve long-horizon discovery behavior.
+Larger models often sustain prolonged exploration of alternative hypotheses, leading to delayed commitment and reduced adaptability under fixed experimental budgets.
+In contrast, small and mid-sized models often exhibit more stable cumulative progress, earlier concentration on informative regions, and more reliable stopping behavior.
+These observations show that baseline LLM solvers relying on exhaustive exploration without principled stopping often exhaust their search budgets (Fig.
+1).
+In contrast, SelfAI-enabled solvers achieve substantially higher hit rates while terminating exploration significantly earlier.
+Notably, final performance is often near saturation across methods, with the best results differing only marginally.
+Nevertheless, significant differences exist in performance and diversity metrics, revealing substantial differences in identifying high-quality configurations and effectiveness in avoiding redundant experiments.
+Figure 3: The top three rows present the Score of all solvers across different tasks, which assess the best stopping performance of SelfAI and baselines using different LLMs.
+The bottom three rows present the Diverse Metrics ($\text{AUP}_{D}$) of all solvers across different tasks, evaluating the diversity of solution trajectories.
 Figure 4: Illustration of the optimized trajectory for the SIREN method for image segmentation.
@@
 We show the serialization recommendations provided by LLM through the labeled numbers.
-2.4 Medical Image Analysis
-In medical image analysis, nnU-Net [72] is a landmark framework known for its strong generalization and automated design.
-Despite its adaptive network structure, preprocessing, and training strategies for various segmentation tasks, the exploration of novel architectures persists.
-To address this, nnU-Net-Revisited [73] establishes a comprehensive benchmark including 19 mainstream models (CNN-based [74, 75, 76], Transformer-based [77, 78], and Mamba-based [79, 80]), offering a solid basis for fair evaluation.
-Results are shown in Figs.
-2-3 and Supplementary Tables 1213.
-On the BTCV dataset [81], GPT4-o3-mini achieves the highest Score (0.6875) and Rank 1, demonstrating strong optimization efficacy, while Qwen2.5-7B secures second place, outperforming all larger variants in its family.
-In nnU-Net hyperparameter tuning for the BraTS dataset [82], Qwen2.5-14B attains the top ranking (Score 0.4333), followed closely by GPT4-o3-mini, with Qwen2.5-32B and DeepSeek-R1-70B tying for third.
-Notably, the 7B and 72B models outperformed similar models of medium size (14B–32B), indicating that larger model size does not guarantee better optimization.
-2.5 Imbalanced Node Classification
-Graph neural networks (GNNs), as powerful tools for processing relational data, are playing an increasingly critical role in numerous cutting-edge scientific fields.
-In biomedicine, GNNs have become a core component of protein structure prediction models such as AlphaFold3 [39], enabling precise modeling of spatial interactions between amino acid residues and ushering in a new era in structural biology.
-In drug discovery, they are widely used for molecular property prediction [39, 83, 84], compound-target interaction identification [85, 86], and novel drug generation [87].
-Similarly, GNNs demonstrate exceptional performance in engineering domains, including traffic prediction [88, 89], cybersecurity, and chip design [90, 91].
-Given this broad utility and the sensitivity of GNN performance to hyperparameter settings, we evaluate GraphSAGE on the imbalanced Cora benchmark [92] under the setup of [93].
-Figs.
-2 and 3 represent that the GPT4-o3 series delivered near-optimal results, achieving the fastest discovery of optimal solutions and efficient stopping decision while maintaining high search diversity, underscoring its robust optimization capability in scientific computing scenarios.
-In contrast, Qwen2.5 models, (except 72b), easily delivered consistent but middling results across scales, while larger models generally exhibit late convergence and higher unnecessary exploration, reflecting potential inherent limitations in understanding the structural properties and optimization requirements of specialized domains such as protein interaction networks and molecular graph structures (see Supplementary Table 14 for more results).
-These results indicate that SelfAI can effectively address challenges in GNN hyperparameter optimization.
-The framework’s strong performance on graph learning tasks suggests its potential as a general-purpose optimizer for scientific domains, including protein engineering and drug development.
-2.6 Drug Discovery
-In drug discovery, accurate bioactivity prediction is vital for early-stage virtual screening and compound prioritization, enabling shorter development cycles and lower preclinical costs.
-Following the evaluation practices and modeling standards summarized by Korotcov et al.
-[94], this study employs the Chagas EP20 dataset [95], which measures the activity of compounds against Trypanosoma cruzi, a neglected but medically significant target in tropical disease research.
-The core challenge lies in learning an effective mapping from molecular representations, such as SMILES sequences [96] or graph-based encodings [97] to experimentally measured bioactivity [98].
-We adopt SelfAI to adaptively refine performance, assess search trajectories, and determine principled stopping points for this task.
-Experimental results (see Supplementary Table 15) reveal distinct performance patterns among solvers in hyperparameter optimization.
-Qwen2.5-7B and 14B secure first and second place, respectively, both achieving high Scores with minimal search diversity.
-GPT-4o ranks third, while DeepSeek-R1-7B and 70B tie for fourth.
-A key finding is the clear performance drop at larger scales; Qwen2.5-32B, Qwen2.5-72B, and DeepSeek-R1-32B rank in the bottom half, demonstrating that increased parameters do not guarantee better optimization.
-In contrast, traditional optimization methods (such as GS and BS) ranked poorly in this task, underscoring the clear advantage of large language model-based intelligent optimizers in bioactivity prediction.
-These results collectively demonstrate that in high-dimensional, noisy biochemical scenarios, language model-driven optimizers offer efficient convergence and a superior balance between exploration and exploitation.
-Finally, the ranking heatmap (Fig.
-7) and average performance analysis (Supplementary Table 3) reveal several consistent trends.
-GPT-4o-mini emerges as the most consistently effective solver, strong inductive reasoning, rapid adaptation to emerging evidence, and stable progression across diverse scientific domains.
-Mid-sized models such as Qwen2.5-7B and Qwen2.5-14B display competitive and robust performance relative to their scale, often exhibiting disciplined evidence-based refinement during exploration.
-In contrast, larger models (Qwen2.5-72B and DeepSeek-R1-70B) exhibit higher variance: they perform well on specific tasks but frequently show delayed breakthroughs, unstable trajectories, and inconsistent cumulative progress.
-These patterns suggest that larger models tend to spend more time exploring alternative possibilities rather than leveraging early evidence, resulting in delayed commitment and reduced adaptability during the search process.
 3 Discussion
-The emergence of the autonomous scientific discovery system (ASDS) represents a paradigm shift in scientific research, with the potential to accelerate discovery by delegating structured experimental processes to artificial intelligence.
-Yet most existing agentic ML frameworks and benchmark-driven ASDS, such as MLE-bench [99], emphasize an agent’s ability to generate and execute candidate solutions, offering limited support for adaptive trajectory management, such as reasoning about when exploration should continue, how search strategies should evolve, or why certain regions of the space warrant further investigation.
-SelfAI addresses this gap by integrating structured intent interpretation, multi-agent reasoning, and principled stopping criteria into a general reasoning framework.
-Across diverse domains, this design enables the system to allocate computational effort where it is most informative, rather than exhaustively enumerating the search space.
-Under this goal, SelfAI is meticulously designed as a general-purpose platform capable of performing high-level reasoning across diverse tasks, thereby accelerating the experimental search process.
-Evaluations across a comprehensive benchmark (12 tasks across 6 different domains) uncover that SelfAI strategically allocates limited computational resources to the most promising performance spaces or research directions and terminates unproductive trajectories, exhibiting human-like cognitive flexibility in dynamically adjusting the search trajectory.
-This optimal stopping behavior effectively alleviates the most persistent inefficiencies in current ASDS.
-The newly proposed Score and $\text{AUP}_{D}$ metrics are crucial for quantifying this advantage, revealing SelfAI’s remarkable balance between rapidly finding discovery efficiency (Score) and maintaining healthy exploration diversity ($\text{AUP}_{D}$).
-These metrics further illuminate the limitations of traditional methods like grid search and Bayesian optimization, which often lack the adaptive reasoning required to escape local optima or terminate unfruitful trials.
-Our findings also suggest that in scientific domains, effective scientific discovery demands specialized reasoning capabilities beyond mere model scale.
-Although large-scale models like GPT-4o and DeepSeek-R1-70B occasionally excel in broad exploration, their performance is frequently surpassed by more compact, purpose-driven models such as Qwen2.5-7B and DeepSeek-R1-7B, which demonstrate superior stopping efficiency and search consistency.
-This indicates that for navigating complex non-convex optimization environments, the structured reasoning and hypothesis generation implemented in SelfAI’s cognitive agent may play a more decisive role than model scale alone.
-The framework demonstrates that purpose-driven architectural design can effectively harness the capabilities of moderately-sized models to deliver robust scientific discovery performance.
-The implications of this framework extend far beyond accelerated experimental search and improved benchmark performance.
-First, the architecture of SelfAI embodies a practical blueprint for human-AI symbiosis in science.
-The User Agent acts as a seamless translator, allowing human researchers to interact with the system at the level of scientific intent and high-level questions, rather than through low-level code or configuration files.
-This design formally integrates human expertise as a guiding input, creating a collaborative loop where the AI manages the scale and complexity of the search while the human provides strategic direction and domain insight.
-This addresses a critical shortcoming of "black-box" automation by fostering reproducibility and trust, as every action of the AI is traceable to a human-defined objective.
-Furthermore, this closed-loop process allows researchers to shift from using AI for scientific research to exploring how to conduct scientific research optimally and to investigate efficient search strategies and problem-related endpoints under controlled conditions.
-SelfAI has the potential to evolve from a tool into a collaborative research partner, enhancing the cognitive and experimental capabilities of human scientists.
-This evolution would be driven by several key advancements.
-For instance, the User Agent could be enhanced to incorporate autonomous code generation, thereby broadening the range of experimental configurations without requiring explicit user pre-definition.
-Furthermore, while the current iteration of SelfAI relies on the static knowledge embedded within its underlying LLM (Large Language Model), future versions could leverage advanced memory management and RAG approaches to alleviate context length limitations and dynamically integrate a richer, more up-to-date body of domain-specific knowledge.
-Collectively, these envisioned capabilities underscore SelfAI’s foundational role in shaping a more adaptive, collaborative, and intelligent paradigm for autonomous scientific discovery.
-4 The Design of SelfAI for Scientific Discovery
-SelfAI is a general scientific discovery system designed to automate and accelerate the end-to-end scientific discovery process.
-Its core function is to transform users’ high-level research intentions (hyperparameter optimization, novel algorithm design, model architecture validation, or ablation studies) into structured, executable workflows.
-The system operates in an integrated loop, comprising three core agents: a user agent, a cognitive agent, and an experiment manager, which together constitute the overall experimental workflow in Fig.
+Autonomous scientific discovery systems (ASDS) have been widely proposed as a means to accelerate research by delegating large-scale experimentation to artificial intelligence.
+However, most existing agent frameworks and ASDS benchmarks primarily emphasize candidate generation and execution, offering limited support for managing long-horizon discovery processes.
+In particular, they provide little explicit reasoning about how search strategies should evolve over time, when exploration should be terminated, or how experimental effort should be redistributed as evidence accumulates.
+This work addresses these limitations by reframing scientific discovery as a trajectory-level decision-making problem.
+By integrating structured intent interpretation, trajectory-aware reasoning, and adaptive stopping within a unified framework, SelfAI enables explicit reasoning over accumulated experimental evidence.
+Across diverse benchmarks, SelfAI consistently reallocates computational resources toward informative regions of the search space and terminates exploration under diminishing returns, reshaping discovery dynamics beyond what is achievable through endpoint-focused optimization alone.
+The newly proposed trajectory-aware evaluation metrics are crucial for quantifying this advantage.
+Metrics such as Score and $\text{AUP}_{D}$ reveal substantial differences in discovery efficiency, redundancy, and stopping behavior that are obscured by conventional endpoint-based evaluation.
+While many solvers ultimately converge to similar final performance, trajectory-level metrics expose how quickly high-quality configurations are identified and how effectively redundant experimentation is avoided.
+These measurements highlight structural limitations of classical optimization methods, including grid search and Bayesian optimization, which lack adaptive reasoning mechanisms for long-horizon exploration and principled termination.
+SelfAI’s trajectory-aware reasoning enables long-term scientific exploration by integrating experimental history.
+While our benchmark results show that LLM-based agents generally exhibit broad exploration behavior and consistent final performance, they can also fall into the trap of excessively long hypothesis exploration times, model execution delays, and greater performance fluctuations.
+Although some small- to medium-sized models, guided by structured reasoning and adaptive stopping, demonstrate more stable trajectory organization and earlier convergence, optimal stopping decisions can still fail in highly complex environments.
+Premature stopping can still lead to the inability to identify the global optimum, while over-exploration can occur under extreme uncertainty (see Supplementary Sect.
+B.6).
+To enhance the long-term scientific exploration capabilities of LLM agents, dynamic knowledge integration and contextual learning capabilities can be integrated to overcome the inherent knowledge bottlenecks of LLMs.
+Beyond benchmark performance, the proposed framework has broader implications for human-AI collaboration in scientific research.
+By separating high-level scientific intent from low-level experimental execution, SelfAI provides a structured interface through which human expertise can guide large-scale automated exploration.
+This design promotes transparency and reproducibility by ensuring that exploration policies remain anchored to explicit research objectives.
+More broadly, it enables a shift from using AI solely as an optimization tool toward studying scientific discovery itself as an object of inquiry—allowing researchers to analyze, compare, and refine discovery strategies under controlled conditions.
+Finally, LLM-based agents enable explicit modeling of long-horizon scientific discovery, shifting discovery systems from task-specific performance optimization toward continuous discovery processes for broader scientific domains.
+Considering a wide range of scientific fields, including physical and materials sciences, there is broad interest but limited expertise in general-purpose LLM agents.
+SelfAI can interleave LLM agents from different application domains to support the design of entirely new systems.
+Moreover, this framework can be integrated with automated toolchains, such as code generation and more powerful inference models, to enable smarter and more adaptive discovery workflows.
+Overall, this work opens new opportunities for scientific discovery by establishing a pathway toward continuous, long-horizon exploration.
+4 Materials and Methods
+4.1 Scientific discovery benchmark
+Existing machine learning benchmarks, such as MLE-bench [chan2024mle], are primarily designed to accelerate the engineering of learning systems by optimizing performance under constrained computational budgets.
+In contrast, we introduce a scientific discovery benchmark specifically tailored to long-horizon discovery, formulated as a trajectory-driven decision-making process.
+Rather than emphasizing final performance, the benchmark requires solvers to engage in scientific reasoning by analyzing experimental outcomes, interpreting accumulated observations, and adaptively revising search strategies to uncover functional relationships, structural patterns, or optimal experimental configurations.
+The benchmark comprises 12 tasks spanning six scientific categories, including scientific computing, machine learning, computer vision, medical image analysis, and drug discovery (Table 1).
+All benchmark data are collected from real experimental runs, with several tasks corresponding to configurations reported in published literature.
+The tasks cover both discrete and continuous search spaces and range from low- to high-dimensional settings.
+To simulate realistic user-system interaction, each task provides background information to the LLM to support intent interpretation.
+Most datasets are generated through systematic grid-based exploration, while LCBench and the Chagas EP20 drug discovery dataset are constructed from Bayesian optimization trajectories.
+LCBench is a widely used AutoML benchmark [zimmer2021auto].
+All tasks are organized as YAML specifications containing metadata, task definitions, concise problem descriptions, search spaces, and trial records.
+The benchmark will be publicly released at https://github.com/XiaoXiao-Woo/SelfAI.
+4.2 Compared methods and evaluation setup
+We evaluate SelfAI across all 12 benchmark tasks and compare it with classical optimization methods, including grid search and the Tree Parzen Estimator (TPE) optimizer [optuna], referred to as BS.
+To assess robustness across model families, we instantiate SelfAI with different language-model backbones, including OpenAI-o3 [openai2023gpt], Llama3.3 [llama], Qwen2.5 [bai2023qwen], and DeepSeek-R1 [guo2025deepseek].
+All methods are evaluated under identical computational settings.
+Reasoning traces are retained for analysis and reproducibility.
+For all trials, the random seed is fixed, and the temperature is set to zero.
+The detailed settings of the benchmark are available in the Supplementary Materials.
+Table 1: List of tasks in SelfAI with different hyperparameters for multiple tasks and datasets.
+Category Method Task Dim Count Ref.
+Scientific Computing Tensor Network Image Completion 3 64 [tw]
+Machine Learning Random Forest Pricing Prediction 5 162 [boston]
+LSTM Sentiment Analysis 2 20 [sentiment_analysis]
+GraphSAGE Node Classification 22 25 [yan2023unreal]
+Computer Vision SIREN Image Denoising 2 25 [sitzmann2020implicit]
+SIREN Image Segmentation 2 25 [sitzmann2020implicit]
+ResNet Image Classification 4 9 [resnet]
+MAE Image Classification 2 20 [mae]
+FashionMnist-NN Image Classification 5 2000 [zimmer2021auto]
+Medical Image Analysis nnUnet BraTS [brats] 3 18 [nnunet]
+nnUnet-revisited BTCV [btcv] 5 19 [nnunet_revisit]
+Drug Discovery DNN Bioactivity Prediction 4 30 [korotcov2017comparison]
+4.3 SelfAI system
+The SelfAI system operates in an integrated loop, comprising three core agents: User Agent, Cognitive Agent, and Experiment manager, which together constitute the overall experimental pipeline in Fig.
 1.
-4.1 User Agent
-The User Agent, as the user interface of SelfAI, is designed to be interactive, user-controllable, and adaptable to evolving research needs.
-It interprets high-level scientific intent and translates natural-language objectives into structured, machine-readable experimental configurations.
-Through iterative clarification, it helps researchers specify objectives, constraints, and search spaces, thereby establishing the experimental environment in which the Cognitive Agent operates.
-For example, requests such as “achieve state-of-the-art accuracy on CIFAR-10 using a CNN” or “identify the most influential method for image classification” are reformulated into precise experimental specifications (see Supplementary Material A.1).
-Unlike static intent-generation prompts, the User Agent supports continuous intervention: researchers may pause execution, inspect reasoning traces, adjust constraints, or redesign experimental protocols at any stage without restarting the workflow.
-4.2 Cognitive Agent
-SelfAI’s cognitive agent is central to accelerating scientific discovery.
-After receiving structured experimental configurations from the user agent and accumulated experimental history from the experiment manager, the agent synthesizes all available information, analyzes experimental trajectories, evaluates observed performance trends, and identifies promising regions in the search space.
-This achieves a delicate balance between developing these high-potential regions and exploring areas of greater uncertainty, ultimately deriving the next optimal search strategy.
-A significant feature of our cognitive agent is its introduction of principled stopping criteria, addressing a key limitation of traditional AI systems, which often continue exploring under conditions of diminishing returns.
-By conducting evidence-based evaluations of research progress, the agent can determine when further experiments on a particular hypothesis are unlikely to yield significant new insights, thereby reallocating resources to more promising research directions (see Supplementary Material A.2 for more details).
-This capability improves resource utilization efficiency in the automated discovery process.
-4.3 Experiment Manager
-The Experiment Manager executes the experimental plans proposed by the Cognitive Agent, each of which has already been structured as executable training jobs, specifying model architectures, datasets, and hyperparameters.
-For every job, the manager initializes these specified configurations, launches the training and evaluation pipeline, and records all associated outputs.
-It tracks runtime status, handles execution failures, and captures performance metrics as well as relevant logs and metadata.
-These records constitute a continuously updated structured knowledge base, providing a reliable foundation for the cognitive agent to analyze search progress, identify promising regions of the search space, and decide when to stop exploration.
-Comprehensive execution procedures and implementation details are provided in Supplementary Material A.3.
-Beyond executing trials, the Experiment Manager also manages the computational environment in which scientific experiments are carried out.
-Through the Experiment Manager, resources are allocated reliably, execution workloads are scheduled efficiently, and checkpointing and recovery mechanisms are coordinated seamlessly.
-Anomalies or failed runs are surfaced to the Cognitive Agent as feedback for complex adaptive reasoning.
-As a result, the overall research process is conducted within an efficient, reliable, and reasoning-driven environment.
+It consists of three components: a User Agent for intent formalization, a Cognitive Agent for iterative decision-making, and an Experiment Manager for experiment orchestration and management.
+Before illustrating the details of SelfAI, we first situate SelfAI within the landscape of existing AI-assisted research frameworks by examining their functional coverage (Table 2).
 Table 2: Comparison of SelfAI with related AI research frameworks and benchmarks across system-level, agent-specific, and task-specific capabilities.
-Functions Ours Code LLaMA [53] MLGym [48] AI Co-Scientist [45] AIRA [24] MLAgentBench [49] Optuna [61]
+Functions Ours Code LLaMA [kochnev2025optuna] MLGym [MLgym] AI Co-Scientist [ai_co_scientist] AIRA [aira] MLAgentBench [mlagentbench] Optuna [optuna]
 Interactive Research
@@
 (*) denotes basic support.
-4.4 Comparison of AI research systems
-We contextualize SelfAI’s capabilities within the current landscape of AI-assisted research tools.
-Table 1 presents a systematic comparison against representative frameworks across multiple dimensions of scientific functionality.
-Existing ASDSs often specialize in code- or idea-level hypothesis generation under fully autonomous pipelines that primarily optimize medal rates within 24 hours.
-In contrast, SelfAI integrates intent interpretation, iterative scientific reasoning, and adaptive trajectory optimization into a unified discovery workflow.
-These capabilities allow the system to refine strategic planning based on accumulated experimental evidence, rather than executing predetermined and static generalization procedures.
-Overall, these advantages position SelfAI as a general-purpose framework capable of supporting efficient, reliable, and reasoning-driven scientific workflows that extend the scope of existing agent or hyperparameter optimization benchmarks.
-5 Data availability
-All datasets used in this study are available for download at https://github.com/XiaoXiao-Woo/SelfAI.
-6 Code availability
-The SelfAI repository is available as Supplementary Software. Updated versions can be found at https://github.com/XiaoXiao-Woo/SelfAI.
+4.3.1 User agent
+As illustrated in Fig.
+1b, User Agent serves as the user interface through which high-level research intent is translated into structured, machine-readable experimental specifications.
+Its primary function is to formalize natural-language objectives provided by users into a standardized configuration that defines the experimental task for the SelfAI system.
+For example, user requests such as “Design high-performance deep learning for image classification task on the ImageNet dataset” or “Identify the most influential method for image classification” are reformulated into precise experimental specifications.
+Specifically, User Agent maps user-provided descriptions of research goals, constraints, and preferences into a task-specific YAML configuration.
+This configuration explicitly specifies basic ideas, optimization objectives, the experimental search space, trial budget, and historical trial data (as shown in Supplementary Sect.
+A.1).
+The resulting YAML file serves as the sole interface between user intent and downstream system components, including Cognitive Agent and Experiment Manager.
+Consequently, User Agent encodes the user’s research intent into a fixed configuration schema, which is implemented via a predefined prompt template.
+All interactions are restricted to the generation and refinement of the structured configuration, and users do not directly manipulate experimental parameters or outcomes once the configuration is finalized.
+4.3.2 Cognitive agent
+Cognitive Agent performs iterative decision-making within the SelfAI framework by operating on structured experimental inputs and conducting trajectory-level analysis across multiple iterations.
+To support long-horizon optimization, Cognitive Agent operates in three functional stages: hypothesis generation, strategic planning, and stopping judgment.
+Hypothesis generation identifies promising regions of the search space based on accumulated experimental outcomes.
+Strategic planning translates these hypotheses into concrete experimental proposals that balance refinement of high-performing configurations with exploration of under-sampled or uncertain regions.
+Stopping judgment evaluates accumulated experimental evidence to determine whether continued experimentation is warranted under the current exploration strategy.
+At each iteration, Cognitive Agent receives a task-specific experimental configuration from User Agent, along with accumulated experimental histories and performance metrics from Experiment Manager.
+These inputs are provided in a structured format and define the current state of the search process.
+Using the accumulated evidence, Cognitive Agent evaluates experimental trajectories, assesses observed performance trends, and estimates coverage of the explored search space.
+Based on this analysis, it generates candidate experimental configurations for subsequent exploration.
+In addition to proposing new experiments, the Cognitive Agent produces a stopping decision based on trajectory-level evaluation of completed trials, including comparison with the initial configuration, assessment of diminishing returns in recent performance improvements, and estimation of the potential value of remaining unexplored regions.
+The outputs of Cognitive Agent consist of candidate experimental configurations for the next iteration and a stopping signal indicating whether the current exploration process should be terminated.
+Detailed implementation, including the prompt templates used for hypothesis generation, strategic planning, and stopping judgment, is provided in the Supplementary Materials A.2.
+4.3.3 Experiment manager
+Experiment Manager is responsible for experiment orchestration and recovery, including resource management, task allocation, and progress tracking (Fig.
+1b).
+These capabilities enable efficient coordination of multi-instance parallel optimization, maximize resource utilization, and enhance training robustness.
+1) Resource Management.
+Experiment Manager monitors resource consumption and dynamically allocates available GPU, TPU, and memory resources.
+This granular allocation optimizes workload distribution across computing units and ensures stable execution of all trials.
+2) Fault recovery and Checkpoint Reconnection.
+In case of system interruptions or suboptimal model performance, Experiment Manager reports failures to Cognitive agent.
+Experiment Manager performs preliminary diagnostics, identifies potential issues, adjusts training parameters, and resumes training from the latest checkpoint.
+3) Multi-Instance Parallel Optimization.
+SelfAI instantiates each user program to run across diverse physical environments, independent of the target program framework.
+Experiment Manager coordinates multi-instance parallel training, synchronizes execution, and concurrently tests various configurations, thereby shortening overall training time and improving generalization across datasets and model parameters.
+For each parallel experiment, Experiment Manager identifies and supplies the necessary runtime parameters, ensuring experiments are conducted under the same environment optimized by Cognitive agent.
+4.4 Evaluation for reasoning trajectories
+While the reasoning process of LLMs in problem-solving often generates a diverse range of discrete insights and multiple potential chains of thought, this diversity, though valuable for exploration and exploitation, can pose a challenge to coherent reasoning evaluation across various experimentation and discovery phases.
+We propose a systematic evaluation metric that captures both the diversity of reasoning perspectives and the overall coherence of the reasoning process, ensuring a more robust and comprehensive assessment of reasoning capabilities.
+Optimal Stopping Criteria In this work, we collected the best value point and the stop point from trials.
+Based on Optimal Stopping Criteria [hill2009knowing], we can define the following measure formulas,
+$\text{Gain}=\frac{1}{N}\sum^{N-1}_{i=0}\frac{v^{*}_{i}-v_{i,\min}}{v_{i,\max}-v_{i,\min}}$ (1)
+$t_{\text{best}}=\frac{1}{N}\sum^{N-1}_{i=0}t^{\text{best}}_{i}=\frac{1}{N}\sum^{N-1}_{i=0}\frac{m_{i}}{M_{i}}$ (2)
+$t_{\text{stop}}=\frac{1}{N}\sum^{N-1}_{i=0}t^{\text{stop}}_{i}=\frac{1}{N}\sum^{N-1}_{i=0}\frac{n_{i}}{M_{i}}$ (3)
+and where $N$ is the number of tasks.
+For the $i$-th task, $M_{i}$ is the number of completed trials.
+$m_{i}$ is the best value point index.
+$t_{i}^{\text{best}}$ is the cost of the best value.
+In addition, we set the stop point index, $n_{i}$, and $t_{i}^{\text{stop}}$ is better when $n_{i}$ is closer to the best value point index.
+$S_{i}$ means the binarized value.
+To obtain a comprehensive measure, we combine the last three measures, i.e., Rel, $t_{\text{best}}$, and $t_{\text{stop}}$, where we utilize Rel in underestimated penalty, then $t_{\text{best}}$ and $t_{\text{stop}}$ are the time penalty ($P_{\text{best}}$ and $P_{\text{stop}}$).
+Thus, the total penalty is
+$P_{\text{total}}$ $=\frac{t_{i}^{\text{stop}}+t_{i}^{\text{best}}}{2}$ (4)
+Finally, the score is denoted as
+$\text{Score}=\frac{1}{N}\sum^{N-1}_{i=0}\text{Gain}\cdot(1-P_{total})$ (5)
+Best Approximation/Candidate In [MLgym], performance profiles and the AUP aim to measure available rates across m tasks, where all performance metrics are threshold $\tau$, performance profiles ($\rho_{\texttt{m}}(\tau)$) are computed as thresholds in all metrics (sorted by ascending) in the current task.
+$\text{AUP}_{\texttt{m}}=\int^{\tau_{\max}}_{1}\rho_{\texttt{m}}(\tau)d\tau$ (6)
+It is noted that the above performance profile and $\text{AUP}_{\texttt{m}}$ score cannot measure the diversity of reasoning. Therefore, we rewrite the performance profile and AUP score:
+First, the performance profile is defined in all completed trials $M_{i}$ in $i$-th task and the overall search space $\mathcal{H}$, as follows
+$r_{i}=\begin{cases}\frac{\max{\{v_{k}:\,k\in\mathcal{H}\}}}{v_{i}},\,\text{ascend}\\ \\ \frac{v_{i}}{\min{\{v_{k}:\,k\in\mathcal{H}\}}},\,\text{descend}\end{cases}$ (7)
+where ascend/descend denotes that the value is larger/smaller and the performance is better.
+For all trials, $\tau$ is the set of all obtained $r_{i}$ values.
+Then, we consider all completed trials $M_{i}$ in $i$-th task,
+$\rho_{i}(\tau)=\begin{cases}|\{k\in M_{i}:r_{k}>=\tau\}|,&\text{ascend}\\ |\{k\in M_{i}:r_{k}<=\tau\}|,&\text{descend}\end{cases}$ (8)
+which captures how many evaluated configurations exceed a given performance threshold $\tau$. $\rho_{i}(\tau)$ is the cumulative distribution curve of the trajectory.
+The area term aggregates the overall concentration of strong configurations along the performance axis:
+$A=\frac{1}{N}\sum^{N-1}_{i=0}\int_{\tau_{\min}}^{\tau_{\max}}\rho_{i}(\tau)d\tau.$ (9)
+To capture the temporal asymmetry of discovery, we compute the centroid
+$G=\frac{1}{A}\int_{\tau_{\min}}^{\tau_{\max}}x\cdot\rho_{i}(\tau)\,d\tau,$ (10)
+and define the skewness
+$S=\int_{\tau_{\min}}^{\tau_{\max}}\left(\frac{x}{G}\right)^{3}\rho(x)\,d\tau.$ (11)
+Since $S$ may be unbounded and may take both positive (left-skewed) and negative (right-skewed) values, we normalize it via a reference skewness value $S_{\text{base}}$ from the GS method and a smooth monotonic mapping:
+$S’=1-\frac{S-S_{\text{base}}}{S_{\text{base}}},$ (12)
+$S^{\prime}=\frac{\tanh(S)+1}{2},S^{\prime}\in(0,1).$ (13)
+Finally, the Area Under the Performance-Diversity curve ($\text{AUP}_{D}$) is defined as
+$\text{AUP}_{D}=A/S^{\prime},$ (14)
+where trajectories that exhibit earlier concentration of high-performing configurations obtain larger and thus smaller $\text{AUP}_{D}$ values, whereas trajectories that concentrate improvements later yield smaller and therefore larger values.
 Appendix A LLM Agents in SelfAI System for Scientific Discovery
@@
 It integrates three core agents: the User Agent, the Cognitive Agent, and the Experiment Manager, and provides an integrated toolkit for autonomous model training, experiment orchestration, and advanced reasoning, enabling a scalable and adaptive workflow.
-A.1 User Agent: Idea Interaction and Experiment Configuration
-In autonomous research systems, the User Agent is typically powered by a general-purpose LLM and acts as the primary interface between human researchers and the system.
-These systems build an “idea-to-experiment” autonomous research process where user access to detailed experimental workflows is limited.
-Such approaches can streamline research execution but may also restrict researchers’ flexibility in iteratively shaping and adapting experiments.
-By contrast, our User Agent unifies idea generation and experiment configuration into a continuous, interactive process, empowering researchers to refine initial concepts, explore alternative approaches, and adjust experimental parameters at any stage.
-During the idea interaction process, the User Agent understands the task background, dissects method details, and defines precise experimental objectives across diverse scientific domains.
-At this stage, the User Agent can help users pinpoint innovative research directions and generate field-relevant ideas.
-Once an idea emerges, the configuration interaction phase leverages prompt templates to clarify user objectives and formulate strategic experimental plans, including initial hypotheses and search configurations.
-This process converts high-level manual queries into standardized and reproducible experimental configurations, as illustrated in the following examples.
-We insert the specified template into the prompt and trigger the prompt to fill out the specified configuration file.
-Consequently, the User Agent can transform ideas into system and user information.
-The system information provides the Cognitive Agent with essential content, including role-playing instructions and task-specific knowledge (such as task benchmarking, method evaluation, comparative analysis, ablation studies, and experimental designs).
-This information equips the Cognitive Agent to interpret high-level research objectives accurately and to execute relevant reasoning and planning steps effectively.
-In parallel, the user information focuses on the practical aspects of experimental planning by supplying detailed specifications and keywords, such as search space definitions, the number of trials, optimization criteria, and resource constraints, which are critical for configuring and running experiments.
-Together, these two types of information enable seamless coordination between conceptual ideation and concrete execution, supporting a workflow in which researchers can iteratively refine their experiments and monitor progress in real time.
+A.1 Prompt of User Agent: Idea Interaction and Experiment Configuration
 User Agent Prompt for Configuration Interaction System: User: {{SUMMARIZED_CONTENT}} Please fill out the content following the YAML format: 1# Template for Configuration Interaction 2# to convert Idea Interaction into YAML format 3- role: system 4 content: 5 model: \$modelName 6 description: You are a \$role specializing in 7 studying \$taskName.
@@
 17- role: user 18 content: 19 max_trials: \$maxTrials 20 trials: []
-A.2 Cognitive Agent
-Figure 5: Illustration of the Cognitive Agent.
-The overall reasoning process involves several key steps: Hypothesis Generation (analysis of the current task and completed trials), Stopping Judgment, and Strategic Planning.
-Strategic Planning develops experimental schemes based on the analyzed hypotheses.
-In our SelfAI framework, the Cognitive Agent continuously ingests and adapts domain knowledge to tackle complex reasoning problems during experiments, much like a human researcher.
-At each stage, the agent first conducts a comprehensive analysis of the task and completed experiments to clarify the purpose and key information of the task and generates preliminary hypotheses.
-Then, it evaluates whether these hypotheses are worth trying through stop judgment to inform strategic planning.
-It then generates specific schemes to probe unexplored regions of the solution space, creating a non-Markovian chain of evolving plans grounded in prior states.
-A.2.1 Reasoning for Scientific Discovery
-LLM agents exhibit remarkable reasoning capabilities, making them well-suited for scientific discovery.
-As illustrated in Fig.
-5, we present an advanced reasoning framework for generating optimal stopping solutions, an approach that can be extended to various scientific discovery tasks.
-A. Trajectory Analysis and Hypothesis Generation
-Specifically, the agent begins with Task 1: Analyze the Current Task, where it meticulously interprets the prompt to establish the task’s core objectives, constraints, and hyperparameters.
+A.2 Reasoning Prompt for Scientific Discovery
+A.2.1 Trajectory Analysis and Hypothesis Generation
+Cognitive agent begins with Task 1: Analyze the Current Task, where it meticulously interprets the prompt to establish the task’s core objectives, constraints, and hyperparameters.
 This initial analysis is used to perform a structured trajectory evaluation, creating a coherent chain of thought that guides all subsequent strategy and exploration.
-This initial analysis is used to perform a structured trajectory evaluation, creating a coherent chain of thought that guides all subsequent reasoning and planning strategies.
 Next, in Task 2: Analysis of Completed Trials, the agent systematically evaluates previous outcomes to generate new hypotheses.
@@
 Step 3: Highlight promising hyperparameter combinations.
-B.
-Best Stopping Judgement As described in Sect.
-LABEL:sec:self-evaluation, we introduce the optimal stopping criterion to guide the prompt design that balances exploration and exploitation.
-During Best Stopping Judgement, the cognitive agent first determines whether the current performance metrics surpass those of the initial configuration and conducts an optimal stopping judgment.
+A.2.2 Stopping Judgement
+We introduce an optimal stopping criterion to guide prompt designs that balance exploration and exploitation.
+During the optimal stopping judgment, Cognitive agent first determines whether the current performance metrics surpass those of the initial configuration and conducts an optimal stopping judgment.
 Building on the prior analysis of completed trials, the agent systematically evaluates stopping conditions to avoid testing all configurations.
@@
 Otherwise, Answer: No with confidence score: {{CONFIDENCE_SOCRE}}.
-C.
-Strategic Planning Following the reasoning and stopping judgment phases shown in Fig.
-5, the agent reviews all executed trials to synthesize key observations on performance trends, notable configurations, and areas of consistent improvement.
-Next, it translates its abstract hypotheses into a concrete experimental plan.
-This strategic process balances two competing objectives: exploiting promising configurations and exploring uncertain regions.
-Exploitation focuses on refining or building upon previously successful configurations, whereas exploration targets under-examined areas that show significant potential for performance gains.
-Each proposed trial is explicitly justified by insights from the preceding reasoning phase, ensuring the search is deliberate and well-founded.
-Finally, each recommendation is formalized as an intended trial and delegated to the Experiment Manager for execution.
-As trials advance and incremental improvements diminish, the Experiment Manager maintains a dynamic record of completed actions and continuously refines the remaining search space.
-Strategic planning directs the optimization trajectory toward convergence, and the system employs an effective stopping strategy to terminate exploration in a rational and efficient manner.
-Notably, the LLM-based agent demonstrates human-like exploratory behavior: it initially focuses on refining hypotheses around high-performing regions and gradually extends attention to areas of greater uncertainty.
-This pattern reflects a synergy of intelligence-driven reasoning and adaptive planning.
+A.2.3 Strategic Planning
+This prompt implements the strategic planning stage of Cognitive Agent.
+Its objective is to generate candidate experimental configurations for the next iteration based on accumulated experimental evidence, while balancing refinement of high-performing configurations and exploration of under-sampled regions.
 Cognitive Prompt for Strategic Planning User: Instructions: Task 1: Review Analysis of All Completed Trials Completed trials: {{COMPLETED_TRIALS}} The following **Search Space** contains **unexplored** trials: {{TRIALS}} Instructions: Task 2: Optimization Recommendation Recommend exactly {{N_JOBS}} promising trials from the provided **Search Space** (include both number and params).
@@
 You MUST provide reasoning for each recommendation.
-A.3 Experiment Manager
-The Experiment Manager is responsible for experiment orchestration and recovery, including resource management, task allocation, and progress tracking.
-These capabilities enable efficient coordination of multi-instance parallel optimization, maximize resource utilization, and enhance training robustness:
-1) Resource Management.
-The Experiment Manager monitors user program resource consumption and dynamically allocates available GPU, TPU, and memory resources.
-This granular allocation optimizes workload distribution across computing units and ensures stable execution of all trials.
-2) Fault recovery and Checkpoint Reconnection.
-In case of system interruptions or suboptimal model performance, the Experiment Manager reports failures to the cognitive agent.
-The Experiment Manager performs preliminary diagnostics, identifies potential issues, adjusts training parameters, and resumes training from the latest checkpoint.
-3) Multi-Instance Parallel Optimization.
-SelfAI instantiates each user program to run across diverse physical environments, independent of the target program framework.
-The Experiment Manager coordinates multi-instance parallel training, synchronizes execution, and concurrently tests various configurations, thereby shortening overall training time and improving generalization across datasets and model parameters.
-For each parallel experiment, the Experiment Manager identifies and supplies necessary runtime parameters, ensuring experiments are conducted under the same environment optimized by the cognitive agent.
-SelfAI maintains detailed logs of all configurations, training runs, and evaluation metrics, and provides targeted optimization suggestions for each completed experiment.
-This intelligent decision-making refines experimental design and reduces redundant search in automated scientific discovery.
-The collaboration between the Experiment Manager and the cognitive agent establishes a self-improving workflow.
-By conducting advanced reasoning and analyzing trial result patterns through hypothesis-driven exploration, they uncover relationships between settings and model performance, steering the search toward optimal configurations.
-This ensures that parallel training is not merely a brute-force process, but an intelligent exploration of the training space that adapts based on real-time outcomes.
-A.4 Evaluation for Reasoning Trajectories
-While the reasoning process of LLMs in problem-solving often generates a diverse range of discrete insights and multiple potential chains of thought, this diversity, while valuable for exploration and exploitation, can pose a challenge to coherent reasoning evaluation across various experimentation and discovery phases.
-We propose a systematic evaluation metric that captures both the diversity of reasoning perspectives and the overall coherence of the reasoning process, ensuring a more robust and comprehensive assessment of reasoning capabilities.
-Optimal Stopping Criteria In this work, we collected the best value point and the stop point from trials. Based on Optimal Stopping Criteria [100, 101], we can define the following measure formulas,
-$\text{Gain}=\frac{1}{N}\sum^{N-1}_{i=0}\frac{v^{*}_{i}-v_{i,\min}}{v_{i,\max}-v_{i,\min}}$ (1)
-$t_{\text{best}}=\frac{1}{N}\sum^{N-1}_{i=0}t^{\text{best}}_{i}=\frac{1}{N}\sum^{N-1}_{i=0}\frac{m_{i}}{M_{i}}$ (2)
-$t_{\text{stop}}=\frac{1}{N}\sum^{N-1}_{i=0}t^{\text{stop}}_{i}=\frac{1}{N}\sum^{N-1}_{i=0}\frac{n_{i}}{M_{i}}$ (3)
-and where $N$ is the number of tasks.
-For the $i$-th task, $M_{i}$ is the number of completed trials.
-$m_{i}$ is the best value point index.
-$t_{i}^{\text{best}}$ is the cost of the best value.
-In addition, we set the stop point index, $n_{i}$, where $t_{i}^{\text{stop}}$ is better when $n_{i}$ is closer to the best value point index.
-$S_{i}$ means that the binarized value.
-To obtain a comprehensive measure, we combine the last three measures, i.e., Rel, $t_{\text{best}}$, and $t_{\text{stop}}$, where we utilize Rel in underestimated penalty, then $t_{\text{best}}$ and $t_{\text{stop}}$ are the time penalty ($P_{\text{best}}$ and $P_{\text{stop}}$).
-Thus, the total penalty is
-$P_{\text{total}}$ $=\frac{t_{i}^{\text{stop}}+t_{i}^{\text{best}}}{2}$ (4)
-Finally, the score is denoted as
-$\text{Score}=\frac{1}{N}\sum^{N-1}_{i=0}\text{Gain}\cdot(1-P_{total})$ (5)
-Best Approximation/Candidate In [48], performance profiles and the AUP aim to measure available rates across m tasks, where all performance metrics are threshold $\tau$, performance profiles ($\rho_{\texttt{m}}(\tau)$) are computed as thresholds in all metrics (sorted by ascending) in current task.
-$\text{AUP}_{\texttt{m}}=\int^{\tau_{\max}}_{1}\rho_{\texttt{m}}(\tau)d\tau$ (6)
-It is noted that the above performance profile and $\text{AUP}_{\texttt{m}}$ score cannot measure the diversity of reasoning. Therefore, we rewrite the performance profile and AUP score:
-First, the performance profile is defined in all completed trials $M_{i}$ in $i$-th task and the overall search space $\mathcal{H}$, as follows
-$r_{i}=\begin{cases}\frac{\max{\{v_{k}:\,k\in\mathcal{H}\}}}{v_{i}},\,\text{ascend}\\ \\ \frac{v_{i}}{\min{\{v_{k}:\,k\in\mathcal{H}\}}},\,\text{descend}\end{cases}$ (7)
-where ascend/descend denotes that the value is larger/smaller, the performance is better. For all trials, $\tau$ is the set of all obtained $r_{i}$ values.
-Then, we consider all completed trials $M_{i}$ in $i$-th task,
-$\rho_{i}(\tau)=\begin{cases}|\{k\in M_{i}:r_{k}>=\tau\}|,&\text{ascend}\\ |\{k\in M_{i}:r_{k}<=\tau\}|,&\text{descend}\end{cases}$ (8)
-which captures how many evaluated configurations exceed a given performance threshold $\tau$. $\rho_{i}(\tau)$ is the cumulative distribution curve of the trajectory.
-The area term aggregates the overall concentration of strong configurations along the performance axis:
-$A=\frac{1}{N}\sum^{N-1}_{i=0}\int_{\tau_{\min}}^{\tau_{\max}}\rho_{i}(\tau)d\tau.$ (9)
-To capture the temporal asymmetry of discovery, we compute the centroid
-$G=\frac{1}{A}\int_{\tau_{\min}}^{\tau_{\max}}x\cdot\rho_{i}(\tau)\,d\tau,$ (10)
-and define the skewness
-$S=\int_{\tau_{\min}}^{\tau_{\max}}\left(\frac{x}{G}\right)^{3}\rho(x)\,d\tau.$ (11)
-Since $S$ may be unbounded and may take both positive (left-skewed) and negative (right-skewed) values, we normalize it via a reference skewness value $S_{\text{base}}$ from the GS method and a smooth monotonic mapping:
-$S’=1-\frac{S-S_{\text{base}}}{S_{\text{base}}},$ (12)
-$S^{\prime}=\frac{\tanh(S)+1}{2},S^{\prime}\in(0,1).$ (13)
-Finally, the Area Under the Performance–Diversity curve ($\text{AUP}_{D}$) is defined as
-$\text{AUP}_{D}=A/S^{\prime},$ (14)
-where trajectories that exhibit earlier concentration of high-performing configurations obtain larger and thus smaller $\text{AUP}_{D}$ value, whereas trajectories that concentrate improvements later yield smaller and therefore larger value.
 Appendix B Details of SelfAI Benchmark
-This section provides the complete technical definitions of the metrics used in the main text, including all normalization procedures, ranking aggregation, and scoring rules.
-These details are omitted from the main text for clarity and conciseness.
-Table 2 summarizes the 12 tasks spanning six scientific categories used to evaluate the reasoning capacity of different solvers.
-For each dataset, we supply background information to the LLMs to simulate the user-intent interpretation process (see Sect.
-A.1).
-Most datasets are generated through systematic grid-based exploration, whereas LCBench and the Chagas EP20 drug-discovery dataset are constructed from Bayesian optimization trajectories.
-LCBench, in particular, is a widely used AutoML benchmark  [71].
+In this section, we summarize the averaged performance across all benchmark tasks.
+Classical baselines, including grid search and TPE-based Bayesian optimization, achieve competitive final objective values but obtain low Scores and high $\text{AUP}_{D}$, reflecting redundant late-stage exploration and delayed stopping (Table 3).
+Naive LLM solvers improve early discovery in some tasks but exhibit substantial variability across domains.
+In contrast, SelfAI-driven solvers based on mid-sized models, such as Qwen2.5-7B and GPT4-o3-mini, consistently achieve the highest Scores with substantially lower $\text{AUP}_{D}$ and earlier stopping times, indicating more focused and efficient discovery trajectories.
+Larger models (e.g., Qwen2.5-72B and Llama3.3-70B) do not necessarily yield improved efficiency, highlighting that trajectory-aware reasoning within SelfAI play a more critical role than model scale.
+For completeness, we discuss representative atypical trajectory behaviors in the Results section, while corresponding examples are provided.
+These cases illustrate exploration behaviors and breakdown patterns observed in SelfAI Benchmark, including premature stopping, limited integration of earlier observations, and sensitivity to minor perturbations, and serve to contextualize the aggregate performance trends reported above.
+B.1 Machine Learning
+Boston house pricing prediction.
+We evaluate SelfAI on the Boston housing price prediction task [boston] using a random forest regression model.
+A total of 162 trials were conducted across a search space defined by five hyperparameters: i.e., $\text{n-estimators}=[100,200,300]$, $\text{max-depth}=[\text{None},10,20]$, $\text{min-samples-split}=[2,5,10]$, min-samples-leaf=[1,2,5] and $\text{max-features}=[\text{``sqrt'', ``log2''}]$).
+As shown in Fig.
+3 and Supplementary Table 4, GPT-4o-mini achieves the highest overall Score (0.9811), ranking first among all evaluated solvers.
+This performance reflects its ability to rapidly identify high-performing configurations while terminating exploration shortly thereafter, yielding superior optimization efficiency.
+Within the DeepSeek-R1 family, the 32B and 7B variants rank second and third, respectively, exhibiting a favorable balance between early convergence (low Best-Time) and moderate exploration diversity among open-source models.
+In contrast, larger models such as DeepSeek-R1-70B and Llama3.3-70B demonstrate sustained and consistent exploration but exhibit less effective stopping behavior, leading to prolonged search trajectories and delayed convergence.
+Notably, nearly all solvers reach an identical and high Best Result (approximately 0.841), indicating that differences in final predictive accuracy are minimal.
+Instead, the primary distinction lies in how efficiently and reliably solutions are discovered, underscoring the importance of trajectory-aware reasoning and adaptive stopping in practical scientific optimization.
+Sentiment analysis.
+We perform experiments for sentiment analysis that focus on identifying opinions, emotions, and attitudes expressed in text.
+All experiments are found in [sentiment_analysis] with standard experimental settings, where pre-trained Word2Vec embeddings [word2vec] are used as input features.
+An LSTM network is then employed to model sentence sequences, converting them into a multi-class classification problem.
+This setup serves to evaluate the agent’s reasoning and optimization capabilities in textual domains.
+As reported in Supplementary Table 5, GPT-4o-mini achieves the highest Score (0.8824), consistently identifying the optimal configuration at an early stage while maintaining a balanced level of exploration.
+In contrast, the DeepSeek-R1 and Qwen2.5 series exhibit substantially lower Scores, frequently showing delayed convergence or failing to reach the global optimum.
+Notably, the largest frontier model, GPT-4o, performs markedly worse (Score 0.2745, rank 11), despite its superior scale.
+This result indicates that model size alone does not guarantee effective iterative hyperparameter reasoning in recurrent neural architectures, and that disciplined trajectory reasoning combined with adaptive stopping plays a more decisive role in efficient optimization within textual learning tasks.
+B.1.1 Scientific Computing
+We selected the tensor decomposition method [tw].
+The method is an important tool for high-dimensional data analysis and is crucial in applications such as data compression [li2025lossless], computational acceleration [berezutskii2025tensor], and multi-modal data fusion [steyaert2023multimodal].
+All LLMs are provided with identical mathematical knowledge about TW decomposition generated by GPT-4o.
+As shown in Supplementary Table 6, Qwen2.5 (7b and 14b) and GPT4-o3-mini perform well and rank highly.
+In contrast, Qwen2.5-72B exhibits reduced performance, suggesting that broader general reasoning alone may be insufficient to navigate the specific trade-offs inherent in tensor decomposition.
+In the DeepSeek series, although the DeepSeek-R1-7b model can’t find the optimal solution, its search strategy was acceptable with a moderate Score, demonstrating its ability to perform mathematical reasoning.
+DeepSeek-R1-70b and GPT-4o, despite excellent search diversity and early stopping, received a mediocre score, reflecting that their exploration strategy was not well aligned with the optimization goal of tensor decomposition, possibly failing to find the optimal performance between data compression and computational efficiency.
+B.2 Computer Vision
+A.
+SIREN We employed SIREN (Sinusoidal Representation Networks) [sitzmann2020implicit] to evaluate our framework on image segmentation and denoising tasks.
+Leveraging sine activations and coordinate-based inputs, SIREN excels at representing high-frequency signals through continuous implicit representations, making it widely used in scientific computing and physics-based problems.
+However, its performance is highly sensitive to hyperparameters (e.g., learning rate and regularization strength, etc.), requiring careful tuning per dataset.
+The unsupervised nature of SIREN further increases the risk of training instability or divergence with improper settings, making it a strong test case for SelfAI’s hyperparameter optimization capability.
+Fig.
+4 illustrates hyperparameter search trajectories for image segmentation using SIREN (see also Supplementary Figs.
+7, 8, and 9), where the surface, smoothed from the original steep, multi-peak data, reveals distinct search behaviors.
+Given the same three initial points, the tree-structured Bayesian optimizer (BS) follows a spiral trajectory that underutilizes promising starting regions and fails to explore broadly, often converging to local minima.
+In contrast, LLM-based optimizers infer trends from evaluated points, incorporate causal reasoning, and explore more broadly to locate the global optimum efficiently.
+They also monitor progress and halt when improvements plateau, a capability absent in traditional methods.
+Quantitative results (Supplementary Tables 7-8) reveal distinct task-dependent performance profiles.
+In segmentation, DeepSeek-R1-7B achieves the highest Score (0.693) and ranks first, followed by Qwen2.5-7B and GPT-4-o3-mini.
+Larger models such as Qwen2.5-72B and Llama-3.3-70B rank only mid-range, indicating that scale alone does not ensure effective optimization.
+In denoising, the landscape shifts: Qwen2.5-72B delivers the best performance (Score 0.761), while DeepSeek-R1-14B and DeepSeek-R1-32B jointly occupy second place.
+Notably, models that perform well in segmentation tasks do not necessarily perform well in denoising tasks, suggesting that the effectiveness of the solver largely depends on the degree of matching between its inference strategy and the task.
+B.
+Image Classification We also benchmark two commonly used supervised learning methods in computer vision: Masked Autoencoder (MAE) [mae] and ResNet [resnet].
+These represent fundamentally different learning paradigms where hyperparameter optimization plays a crucial role in achieving the latest performance.
+Mask Autoencoder (MAE) MAE introduces two key hyperparameters: training strategies (linear detection vs.
+fine-tuning) and masking rate, which are explored over a range: $[0.10,0.90]$ with an interval of 0.1.
+The masking rate directly affects the difficulty of the reconstruction task and the quality of the learning representation, while the training strategy determines how to adapt to the downstream task.
+In our experiments (see Supplementary Table 9 for more details), GPT-4o-mini achieved explicit performance with efficiency, demonstrating the extraordinary ability to identify the best mask configuration and training strategies.
+Qwen2.5-7B achieves early convergence and disciplined stopping.
+The DeepSeek-R1 series shows consistent scaling behavior: performance improves from 7B to 14B to 32B, but slightly drops at 70B.
+In traditional solvers, GS and BS lag significantly behind LLM-based solvers, highlighting the advantages of LLM-enabled discovery.
+Apart from Qwen2.5-14b and DeepSeek-R1-7b, most solvers can find the optimal solution, but significant differences exist among the various methods in terms of search efficiency and convergence speed.
+ResNet Family On the ImageNet ResNet hyperparameter benchmark, LLMs in the 7B-32B range demonstrate superior optimization efficacy.
+The top-performing solvers, such as Qwen2.5-14B (1st), DeepSeek-R1-32B (2nd), and Qwen2.5-7B with DeepSeek-R1-14B (tied 3rd), consistently identify the optimal ResNet configuration within one or two trials, halting immediately with near-perfect sample efficiency.
+In contrast, larger variants from the same model families drop to middle or lower rankings, consuming over half the search budget with notably higher exploration overhead.
+While GPT-series models, as large-scale counterparts, still achieve competitive results, the 7B-32B class exhibits a clear advantage in balancing accuracy and efficiency.
+For the architectural search on ImageNet, we explored key dimensions such as depth and bottleneck design.
+Qwen2.5-14B attains theoretical peak performance, with its 7B and 32B versions completing further exploration efficiently.
+These results reflect an ability to identify ideal network configurations with minimal wasted exploration.
+Large-scale benchmark We evaluate SelfAI on the standard AutoML benchmark [zimmer2021auto] using Bayesian optimization over 2000 rounds of hyperparameter search, covering critical parameters including learning rate, batch size, network depth, and dropout rate (Supplementary Fig.
+5).
+These trials exceed the maximum token limit of GPT-family models, thereby SelfAI does not employ GPT4-o3-mini or GPT4-o3.
 Supplementary Fig.
-6 provides an additional structural analysis of the LCBench search space.
-Accuracy depends on interacting hyperparameters such as weight decay, hidden-layer width, and depth, and the high-performance region is sharply localized.
-This structure highlights why solvers must possess strong reasoning capabilities to efficiently locate these promising configurations.
+5a-e presents that accuracy depends on interacting hyperparameters such as weight decay, hidden-layer width, and depth, and the high-performance region is sharply localized.
+The optimization process using the Tree Parzen Estimator (TPE) optimizer is shown with performance fluctuations that cannot be stopped (Supplementary Fig.
+5f), illustrating highly sparse promising regions.
 Supplementary Fig.
-7 reports solver rankings based on the Score metric.
-These rank summaries allow cross-task comparisons of solver consistency; detailed interpretation of performance trends is provided in the main text.
-In Table 3, we summarize averaged performance across all benchmark tasks.
-Classical baselines such as grid search and TPE-based Bayesian optimization achieve competitive final objective values but receive low Score and high $\text{AUP}_{D}$, reflecting redundant late-stage exploration and delayed stopping.
-Naive LLM solvers improve early discovery but remain unstable across tasks.
-In contrast, SelfAI-driven solvers based on mid-sized models such as Qwen2.5-7b and GPT4-o3-mini attain the highest Scores with substantially lower $\text{AUP}_{D}$ and earlier stopping, indicating more focused and efficient trajectories.
-Larger models (e.g., Qwen2.5-72b, Llama3.3-70b) do not necessarily yield better efficiency, highlighting that structured reasoning and trajectory control within SelfAI are more critical than sheer model scale.
-In addition, as discussed in Sect.
-B.1, failure analyses are also included for completeness.
-These examples illustrate typical breakdowns in reasoning processes across long trajectories, such as premature stopping, limited incorporation of earlier observations, or sensitivity to minor perturbations.
-They inform the broader evaluation of solver stability and support the claims presented in the main text.
-Table 2: List of tasks in SelfAI with different hyperparameters for multiple tasks and datasets.
-Category Method Task Dim Count Ref.
-Scientific Computing Tensor Network Image Completion 3 64 [65]
-Random Forest Regression 5 162 [62]
-Deep Learning LSTM Sentiment Analysis 2 20 [63]
-GraphSAGE Node Classification 22 25 [93]
-Computer Vision SIREN Image Denoising 2 25 [68]
-SIREN Image Segmentation 2 25 [68]
-ResNet Image Classification 4 9 [70]
-MAE Image Classification 2 20 [69]
-FashionMnist-NN Image Classification 5 2000 [71]
-Medical Image Analysis nnUnet BraTS [82] 3 18 [72]
-nnUnet-revisited BTCV [81] 5 19 [73]
-Drug Discovery DNN Bioactivity Prediction 4 30 [94]
+5g presents sharply local high-performance regions are presented arising from interacting hyperparameters.
+We explore these regions in Supplementary Table 11, where leading models (DeepSeek-R1-70B, Qwen2.5-14B, and Llama3.3-70B) efficiently identify near-optimal performance by evaluating only a minimal set of candidate solutions, achieving an order-of-magnitude improvement in stop-time compared to classical methods, as promising regions are identified arising from interacting hyperparameters.
+B.3 Medical Image Analysis
+In medical image analysis, nnU-Net [nnunet] is a landmark framework known for its strong generalization and automated design.
+Despite its adaptive network structure, preprocessing, and training strategies for various segmentation tasks, the exploration of novel architectures persists.
+To address this, nnU-Net-Revisited [nnunet_revisit] establishes a comprehensive benchmark including 19 mainstream models (CNN-based [ronneberger2015u, nnunet, roy2023mednext], Transformer-based [vaswani2017attention, tang2022self], and Mamba-based [gu2023mamba, ma2024u]), offering a solid basis for fair evaluation.
+Results are shown in Figs.
+3 and Supplementary Tables 1213.
+We evaluate SelfAI on this benchmark across multiple datasets, with results summarized in Fig.
+3 and Supplementary Tables 12-13.
+On the BTCV dataset [btcv], SelfAI instantiated with GPT4-o3-mini achieves the highest Score (0.6875) and ranks first, indicating rapid identification of high-quality configurations combined with efficient stopping behavior.
+Notably, Qwen2.5-7B secures second place, outperforming all larger variants within the same model family, suggesting that effective optimization in this setting does not scale monotonically with model size.
+A similar pattern emerges in nnU-Net hyperparameter tuning on the BraTS dataset [brats].
+Qwen2.5-14B attains the top ranking (Score 0.4333), followed closely by GPT4-o3-mini, while Qwen2.5-32B and DeepSeek-R1-70B tie for third.
+Across both datasets, smaller (7B) and very large (72B) models outperform medium-sized counterparts (14B-32B), reinforcing the observation that increased model capacity alone does not guarantee superior optimization performance.
+These results highlight the importance of trajectory-aware decision-making under noisy and high-variance evaluation regimes, which are characteristic of medical image analysis.
+By committing early to promising configurations while avoiding prolonged exploration with diminishing returns, SelfAI enables efficient discovery in settings where excessive experimentation is both computationally costly and practically constrained.
+B.4 Imbalanced Node Classification
+Graph neural networks (GNNs) are a central modeling paradigm for relational data and have become indispensable across a range of scientific domains.
+In biomedicine, GNNs underpin modern protein structure prediction systems, such as AlphaFold3, by enabling explicit modeling of spatial interactions among amino acid residues, thereby advancing structural biology at unprecedented resolution [alphafold3].
+In drug discovery, GNNs are widely applied to molecular property prediction, compound-target interaction modeling, and de novo molecule generation [bongini2021molecular].
+Despite their success, optimizing GNNs on imbalanced datasets remains particularly challenging, due to sparse and asymmetric reward signals, complex relational dependencies, and high-dimensional hyperparameter spaces.
+These characteristics make such tasks especially sensitive to exploration strategy and stopping decisions.
+To investigate whether trajectory-aware reasoning can address these challenges, we evaluate GraphSAGE on the imbalanced Cora benchmark [zhao2021graphsmote] under the setup [yan2023unreal].
+As shown in Fig.
+3, SelfAI instantiated with the GPT4-o3 series consistently achieves near-optimal performance, characterized by rapid identification of optimal configurations, efficient stopping decisions, and sustained exploration diversity.
+This behavior indicates that SelfAI effectively balances exploitation and exploration in structurally complex optimization landscapes.
+In contrast, Qwen2.5 models (with the exception of the 72B variant) exhibit stable but middling performance across scales, while larger models tend to converge later and incur substantial redundant exploration.
+This pattern suggests potential limitations in capturing the structural optimization requirements of relational domains such as molecular graphs and protein interaction networks (see Supplementary Table 14).
+Across different language model backbones, SelfAI consistently demonstrates fast discovery of high-quality configurations while avoiding premature convergence.
+These results provide evidence that trajectory-level reasoning and adaptive stopping generalize beyond continuous parameter optimization, extending effectively to structured, relational learning problems that are common in real-world scientific applications.
+B.5 Drug Discovery
+Drug discovery increasingly benefits from advances in machine learning and data-driven modeling, which serve as versatile frameworks for tasks such as bioactivity prediction, virtual screening, and compound prioritization.
+These approaches have demonstrated the potential to accelerate early-stage discovery, reduce experimental burden, and shorten development timelines [brahmavar2024generating, grisoni2021combining].
+Following the evaluation practices and modeling standards summarized by Korotcov et al.
+[korotcov2017comparison], this study employs the Chagas EP20 dataset [Chagas], which measures the activity of compounds against Trypanosoma cruzi, a neglected but medically significant target in tropical disease research.
+The core challenge lies in learning an effective mapping from molecular representations, such as SMILES sequences [weininger1988smiles], to experimentally measured bioactivity [fliri2005biological].
+We adopt SelfAI to adaptively refine model performance, monitor discovery trajectories, and determine principled stopping criteria for iterative optimization.
+Experimental results (see Supplementary Table 15) reveal distinct performance patterns among solvers in hyperparameter optimization.
+SelfAI and select LLM-based solvers achieve superior Scores by rapidly identifying promising hyperparameter regions and terminating exploration once gains plateau.
+Notably, these results demonstrate efficient optimization under realistic biochemical uncertainty, highlighting trajectory-aware optimization as a practical computational enabler for scalable and resource-efficient bioactivity modeling in drug discovery pipelines.
 Figure 5: The relationship between different parameter settings and accuracy.
-a A hexbin plot showing the joint density of learning rate (log-scaled) and accuracy.
-b Box plots illustrating the distribution of dropout rates used for models of different depths.
-c Mean accuracy and standard error across logarithmically binned batch sizes.
-d Performance comparison of different architectural configurations (layers and units).
-e A scatter plot of weight decay against momentum, colored by accuracy.
-f The optimization trajectory, showing accuracy improvement over successive trials, with a moving average trend line.
+a, A hexbin plot showing the joint density of learning rate (log-scaled) and accuracy.
+b, Box plots illustrating the distribution of dropout rates used for models of different depths.
+c, Mean accuracy and standard error across logarithmically binned batch sizes.
+d, Performance comparison of different architectural configurations (layers and units).
+e, A scatter plot of weight decay against momentum, colored by accuracy.
+f, The optimization trajectory, showing accuracy improvement over successive trials, with a moving average trend line.
 g A correlation matrix quantifying linear relationships between all parameters and the accuracy.
@@
 Refer to caption Refer to caption Refer to caption Refer to caption
-Figure 9: More optimized trajectories of Fig. 8. These LLMs have different suggestions for non-convex hyperparameter optimization.
+Figure 8: More optimized trajectories of Fig. 7.
 DeepSeek-R1-7B
@@
 Figure 9: More optimized trajectories of Fig. 7. These LLMs have different suggestions for non-convex hyperparameter optimization.
-B.1 Failure Cases
-Comparison of Best Result Hit Rate.
-Table 3 compares the experimental results of traditional solvers and LLM-based solvers.
-Early-stopping LLM-ES partially alleviates redundant trials but remains less stable and less efficient than SelfAI-integrated variants.
-In contrast, SelfAI consistently enhances trajectory-level reasoning across all LLM families, reducing unnecessary exploration (lower $\text{AUP}_{D}$), improving stopping behavior (lower Stop-Time), and yielding higher overall Score.
-Mid-sized models such as Qwen2.5-7B and DeepSeek-R1-32B demonstrate particularly favorable exploration and exploitation trade-offs.
-Although LLMs possess advanced reasoning capabilities and can identify optimal results more efficiently, we found that LLM-based solvers still face challenges related to context limitations and computational fragility.
+B.6 Details of limitations
 1) Context limitations.
-The LCBench dataset exceeds the maximum token limit of GPT-family models, which is why SelfAI does not employ GPT4-o3-mini or GPT4-o3.
-In practice, search spaces rarely contain thousands of trials, and this issue can be mitigated in future work.
-For instance, memory management and Retrieval-Augmented Generation (RAG) can help reduce token requirements by introducing additional steps to the experimental recommendation workflow, such as embedding, chunking, and retrieval.
+LCBench contains approximately 2000 trials, constituting a large and highly sparse search space.
+Supplementary Fig.
+5 provides a structural analysis of this landscape.
+Performance depends on interacting hyperparameters such as weight decay, hidden-layer width, and depth, and high-performing regions are sharply localized (Supplementary Fig.
+5a-e).
+In practice, such large and sparse search spaces are common in scientific discovery tasks and can exceed the maximum context length supported by current GPT-family models.
 2) Computational fragility.
-Table 3 compares the best result hit rates across solvers.
+Supplementary Table 3 compares the best result hit rates across solvers.
 The performance of LLM-based solvers is highly sensitive to the reasoning strategy, and early stopping often leads to a decline in hit rate.
@@
 As illustrated in Fig.
-11, computational fragility also hinders reliable performance improvement through parameter tuning, as seen in the non-monotonic performance of the DeepSeek-R1 model series.
-Table 6 summarizes the metrics corresponding to Fig.
-11.
-Fig.
-11 provides a detailed visualization of solver behavior.
-The Qwen2.5 models consistently identify high-performing configurations early in the search, with Qwen2.5-7B reaching near-optimal regions within the first 10–15 trials.
-Their trajectories exhibit stable refinement and limited oscillation, suggesting consistent integration of prior experimental evidence.
-In contrast, the DeepSeek-R1 models show pronounced non-monotonic behavior.
-DeepSeek-R1-7B frequently oscillates between high and low values, reflecting sensitivity to minor variations in its reasoning process.
-DeepSeek-R1-70B achieves early improvements but then terminates prematurely, failing to further refine promising regions.
-Although the GS and BS reach high-performance regions, their trajectories lack the rapid breakthroughs observed in LLM-based solvers.
-Overall, these trajectory comparisons provide visual evidence of differences in exploration stability, breakthrough efficiency, and stopping behavior across solver families, complementing the aggregate metrics reported in the main text.
+10, computational fragility also hinders reliable performance improvement through parameter tuning, as seen in the non-monotonic performance of the DeepSeek-R1 model series.
+3) Insufficient trajectory integration.
+Supplementary Fig.
+10 provides representative trajectory visualizations that highlight differences in evidence integration and stopping behavior.
+The Qwen2.5 models typically identify high-performing configurations early, with Qwen2.5-7B reaching near-optimal regions within the first 10-15 trials, followed by stable refinement with limited oscillation.
+In contrast, DeepSeek-R1 trajectories are more frequently non-monotonic.
+DeepSeek-R1-7B often oscillates between high- and low-performing regions, indicating instability in how accumulated evidence is leveraged, while DeepSeek-R1-70B can terminate prematurely after early improvements, leaving promising regions under-refined.
+Classical baselines (GS and BS) may also reach high-performing regions, but their improvements tend to be slower and less punctuated by rapid breakthroughs.
+Overall, these examples complement the aggregate metrics in the main text by illustrating how limitations can manifest as reduced trajectory stability, delayed breakthroughs, or suboptimal stopping decisions.
 Figure 10: Failure Cases for the DeepSeek-R1 Family in the scientific computing field.
@@
 […]
-GS 0.1100 1.000000 0.780000 1.000000 16.63 13
-BS 0.329545 0.761200 0.477273 0.863636 16.63 11
-Qwen2.5-7b 0.602273 0.487619 0.363636 0.431818 16.63 2
-Qwen2.5-14b 0.454545 0.708497 0.272727 0.818182 16.63 4
-Qwen2.5-32b 0.374388 0.538240 0.613636 0.636364 16.60 9
-Qwen2.5-72b 0.408757 0.536178 0.636364 0.545455 16.60 5
-DeepSeek-r1-7b 0.693182 0.318656 0.227273 0.386364 16.63 1
-DeepSeek-r1-14b 0.408534 0.471791 0.704545 0.477273 16.60 6
-DeepSeek-r1-32b 0.312477 0.198417 1.000000 0.340909 15.77 12
-DeepSeek-r1-70b 0.374555 0.576330 0.590909 0.659091 16.60 8
-Llama3.3-70b 0.380344 0.285865 0.590909 0.295455 10.79 7
-GPT4-o3-mini 0.477273 0.517709 0.454545 0.590909 16.63 3
-GPT4-o3 0.363636 0.671366 0.477273 0.795455 16.63 10
+GS 0.1100 1.0000 0.7800 1.0000 16.63 13
+BS 0.3295 0.7612 0.4773 0.8636 16.63 11
+Qwen2.5-7b 0.6023 0.4876 0.3636 0.4318 16.63 2
+Qwen2.5-14b 0.4545 0.7085 0.2727 0.8182 16.63 4
+Qwen2.5-32b 0.3743 0.5382 0.6136 0.6364 16.60 9
+Qwen2.5-72b 0.4088 0.5362 0.6364 0.5455 16.60 5
+DeepSeek-r1-7b 0.6932 0.3187 0.2273 0.3864 16.63 1
+DeepSeek-r1-14b 0.4085 0.4718 0.7045 0.4773 16.60 6
+DeepSeek-r1-32b 0.3125 0.1984 1.0000 0.3409 15.77 12
+DeepSeek-r1-70b 0.3746 0.5763 0.5909 0.6591 16.60 8
+Llama3.3-70b 0.3803 0.2859 0.5909 0.2955 10.79 7
+GPT4-o3-mini 0.4773 0.5177 0.4545 0.5909 16.63 3
+GPT4-o3 0.3636 0.6714 0.4773 0.7955 16.63 10
 Table 8: Performance comparison of different solvers for the image denoising task based on the SIREN model.
@@
 GPT4-o3 0.6023 0.4786 0.2045 0.5909 24.78 6
-Table 9: Performance comparison of different solvers for Mask AutoEncoders (MAE) [69].
+Table 9: Performance comparison of different solvers for Mask AutoEncoders (MAE) [mae].
 […]
@@
 GPT4-o3 0.5833 0.2192 0.3333 0.5000 21.43
-Table 11: Performance comparison of different solvers for the image classification task on the LCBench dataset reported from [71].
+Table 11: Performance comparison of different solvers for the image classification task on the LCBench dataset reported from [zimmer2021auto].
 […]
@@
 GPT4-o3 - - - - - -
-Table 12: Performance comparison of different solvers for the medical image segmentation task on the BTCV dataset reported from the nnUnet benchmark [73].
+Table 12: Performance comparison of different solvers for the medical image segmentation task on the BTCV dataset reported from the nnUnet benchmark [nnunet_revisit].
 […]
@@
 GPT4-o3 0.4063 0.5081 0.4375 0.7500 85.04 4
-Table 13: Performance comparison of different solvers for the nnUnet model on the BraTS dataset [82].
+Table 13: Performance comparison of different solvers for the nnUnet model on the BraTS dataset [brats].
 […]
@@
 GPT4-o3 0.7727 0.2540 0.1818 0.2727 89.282 1
-Table 15: Performance comparison of different solvers for bioactivity prediction on Chagas EP20 dataset [95]
+Table 15: Performance comparison of different solvers for bioactivity prediction on Chagas EP20 dataset [Chagas]
 […]